### v2

# Probing Einstein-Maxwell-Scalar Black hole via Thin Accretion Disks and Shadows with EHT Observations of M87* and Sgr A*

Abstract
[…]
I Introduction
Black holes (BHs) represent one of the fundamental solutions to Einstein’s field equations and are regarded as robust predictions of General Relativity (GR). While historically debated, numerous observational efforts have been made to confirm their existence. A significant breakthrough occurred in 2015 when the Laser Interferometer Gravitational-Wave Observatory (LIGO) detected gravitational waves (GWs) originating from a binary BH merger 124 . Subsequently, in 2019, the Event Horizon Telescope (EHT) provided the first direct observational evidence of a BH shadow at the center of the galaxy M87 125 ; 126 ; 127 ; 128 ; 129 ; 130 .
Further advancements in EHT observations led to the polarization imaging of M87*, revealing the presence of a magnetic field, which provides insights into jet formation mechanisms 131 ; 132 ; 133 . Additionally, EHT reported the detection of the shadow of SgrA*, the supermassive BH at the center of the Milky Way 134 . These observations offer valuable constraints on the geometry of spacetime in the strong-field regime, particularly near the event horizon, allowing precise measurements of BH mass and spin 135 ; 136 ; 145 ; 146 ; 147 ; 148 ; 149 ; 150 ; 151 ; 152 ; 153 ; 154 ; 155 ; 185 . Collectively, these findings provide a crucial platform for testing GR and exploring potential modifications to gravitational theories in extreme astrophysical environments 137 .
It is well known that GR has demonstrated remarkable accuracy in various tests 138 . However, it faces certain limitations within the cosmological framework. These include challenges in explaining the accelerated expansion of the universe, the presence of initial singularities, the missing mass problem, and the Coincidence Problem. Furthermore, testing GR in the strong-field regime, particularly in the vicinity of BH horizons, remains a significant challenge 139 . In addition, modifications to GR have been explored to extend our understanding beyond its conventional framework, aiming to probe near-horizon physics and the large-scale structure of the universe 140 ; 141 ; 142 ; 180 .
Among various modified gravity models, the Einstein-Maxwell-Scalar (EMS) model serves as a promising candidate. Since the EMS theory allows for hairy BH solutions 143 ; 144 ; 145 , where BHs can possess scalar hair, investigating their properties in the strong-field regime is of significant interest. Previous studies have shown that dilaton BH spacetimes are influenced by the dilaton charge, which not only serves as a new form of hair but also provides a powerful tool for analyzing optical phenomena in such backgrounds. In this work, we consider a charged BH solution within the EMS theory as an extension of the Reissner-Nordström (RN) solution. Unlike dilaton BHs, where the scalar field typically contributes through an effective charge, the BH solution in EMS theory is directly governed by its model parameters, which serve as the primary distinguishing feature. A comprehensive exploration of the physical and optical properties of this solution is therefore essential. Investigating the effects of its parameters on spacetime geometry and optical phenomena will contribute to a deeper understanding of its astrophysical implications and provide a means to differentiate it from other BH solutions.
In this article, we analyzed the optical properties of this BH solution through its shadow and the observable appearance of accretion disk emission in the strong-field regime. We further investigated the impact of EMS BH parameters on optical phenomena and constrained the viable ranges of parameters $\alpha$ and $\beta$ using EHT observational data. Our paper is organized as follows. In Sec.II, we will give the BH solution in EMS theory. In Sec.III, we will discuss the influence of EMS parameters on the optical properties of BH and plot the photon orbits of BH. Furthermore, we investigated the observable effects of radiation emitted by the accretion disk and found that the EMS BH exhibits a significantly reduced radiation intensity compared to the Schwarzschild and RN BHs. Finally, we present our conclusions and provide an outlook in Sec. IV. The paper has been written with the unit system $G=M=c=1$, and metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ).
II The black hole solution in Einstein-Maxwell-scalar theory
The interaction between a scalar field and gravity was first systematically studied by Fisher, who obtained a static and spherically symmetric solution to the Einstein equations in the presence of a massless scalar field 102 . This pioneering work laid the foundation for subsequent investigations into the role of scalar fields in gravitational theories. In particular, the EMS theory has garnered considerable interest due to its profound physical implications and its natural emergence in various theoretical frameworks. Notably, it arises in the context of Kaluza-Klein theories, where extra-dimensional reductions lead to scalar field interactions 103 . Furthermore, it plays a crucial role in supergravity and string theory, where dilaton fields naturally couple to gravity and gauge fields 104 . Additionally, scalar fields are fundamental in cosmology, influencing early universe dynamics, inflationary models, and the late-time acceleration of the universe 105 . These diverse connections underscore the significance of the EMS theory in modern theoretical physics.
Within the framework of effective field theory (EFT), the interaction between a scalar field and an electromagnetic field can be incorporated through specific coupling terms in the Lagrangian. Such couplings naturally arise in the low-energy effective description of string theory, where the scalar field is identified with the dilaton—a massless excitation resulting from the compactification of extra spatial dimensions—while the electromagnetic field originates from gauge fields in the higher-dimensional theory. The dilaton, as a key modulus field, encodes information about the compactification geometry and the associated moduli space, dynamically influencing the strength of gauge interactions in the four-dimensional effective theory.
The coupling between the scalar and electromagnetic fields is an inevitable consequence of dimensional reduction, where the dilaton acts as a dynamical field that rescales the effective gauge couplings. This interaction not only modifies the propagation and self-interactions of electromagnetic waves but also introduces additional curvature terms in the gravitational sector, leading to non-trivial corrections to the Einstein field equations when the scalar field is present. These effects are particularly relevant in scenarios involving strong-field gravity, early-universe cosmology, and potential deviations from classical electrodynamics in high-energy regimes. Further implications include modifications to BH solutions, cosmological inflation models, and the possible emergence of new observational signatures in astrophysical and laboratory settings 169 ; 170 ; 171 ; 172 ; 173 ; 174 ; 175 ; 186 .
Such interaction terms are often introduced to match experimental or astrophysical observations under specific conditions. A typical form of the interaction term is given by
$\mathcal{L}_{\text{int}}=f(\phi)F_{\mu\nu}F^{\mu\nu},$ (1)
where $f(\phi)$ is a function of the scalar field $\phi$, and $F_{\mu\nu}$ represents the electromagnetic field strength tensor. Depending on the functional form of $f(\phi)$, the coupling may be constant, exponential, or take other more complex forms. These interaction terms can significantly influence the propagation of electromagnetic fields, especially in regions with strong scalar field backgrounds, such as near BHs or on cosmological scales.
Furthermore, the inclusion of a scalar field can be interpreted as a modification to standard Einstein gravity. Such frameworks are often categorized as scalar-tensor theories or extended theories of gravity. In these models, the coupling between the scalar field and the electromagnetic field can significantly influence the dynamics of the electromagnetic field and modify the solutions of the gravitational field equations. For example, investigations within the EMS theory have demonstrated that the presence of a scalar field can alter the electromagnetic properties of BHs, giving rise to novel classes of BH solutions, such as those characterized by scalar-electromagnetic coupling.
We consider the EMS theory described by the following action
$S[g_{\mu\nu},\phi]=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}(R-2\nabla_{\mu}\phi\nabla^{\mu}\phi-K(\phi)F_{\mu\nu}F^{\mu\nu}-V(\phi)),$ (2)
where $R$ is the Ricci scalar. The EMS theory describes the real scalar field $\phi$ minimally coupled to GR and non-minimally coupled to Maxwell’s background. $K(\phi)$ is the coupling function between the Maxwell and scalar field, and $V(\phi)$ is the scalar potential. if we set up the determined expressions of $K(\phi)$ and $V(\phi)$ first, then the corresponding BH solutions are determined. For example, if we set $K=1$ and $V=2\lambda$, $\lambda$ is the cosmological constant, then the EMS theory gives the Reissner-Nordstrom-de-Sitter solution. And if we set $K=e^{2\phi}$ and $V=0$, then we can get the dilation BH solution 1 ; 2 . Some important solutions with different $K(\phi)$ and $V(\phi)$ are given by 3 ; 4 ; 5 ; 6 ; 7 ; 8 .
In our research, we consider $V=0$ and the coupling function $K(\phi)$ is
$K(\phi)=\frac{(\alpha^{2}+1)e^{\frac{-2\phi}{\alpha}}}{(\alpha^{2}+1+\beta)e^{\frac{-2\phi(\alpha^{2}+1)}{\alpha}}+\beta\alpha^{2}},$ (3)
with $\alpha$ and $\beta$ as characteristic parameters or coupling constants; it is observed that when $\beta\to\infty$, the influence of the electromagnetic field is negligible. Conversely, in the extremal case where $\beta=0$, it simplifies to the Einstein-Maxwell-Dilaton (EMD) coupling described by $K(\phi)=e^{2\alpha\phi}$. The equations of motion are obtained by varying the action with respect to the metric, dilaton, and Maxwell field, respectively, as follows
$\left\{\begin{array}[]{l}0=\nabla_{\mu}\left[K(\phi)F^{\mu\nu}\right],\\ \Box\phi=\frac{1}{4}\frac{\partial K(\phi)}{\partial\phi}F^{2},\\ R_{\mu\nu}=2\partial_{\mu}\phi\partial_{\nu}\phi+2K(\phi)\left(F_{\mu\sigma}F^{~{}\sigma}_{\nu}-\frac{1}{4}g_{\mu\nu}F^{2}\right).\end{array}\right.$ (4)
Due to the field equations mentioned above, Ref.9 provides a BH solution, which can be described as
$\left\{\begin{split}&ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+C(r)(d\theta^{2}+\sin^{2}\theta d\phi^{2})\\ &f(r)=\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{\frac{1-\alpha^{2}}{1+\alpha^{2}}}+\frac{\beta Q^{2}}{C(r)}\\ &C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}},\end{split}\right.$ (5)
where $b_{1}$ and $b_{2}$ are functions solely of $\alpha$, yielding
$\left\{\begin{split}b_{1}&=\left(1+\sqrt{1-q^{2}(1-\alpha^{2})}\right)M,\\ b_{2}&=\frac{1+\alpha^{2}}{1-\alpha^{2}}\left[1-\sqrt{1-q^{2}(1-\alpha^{2})}\right]M.\end{split}\right.$ (6)
with $q\equiv{Q}/{M}$ the charge-to-mass ratio, and $M$ the mass of the BH.
The location of horizons is defined where $f(r_{\pm})=0$. Specifically, for $\beta=0$, these configurations align with the GMGHS solutions, as referenced in 11 , where $r_{\pm}$ are identified as $b_{1}$ and $b_{2}$, respectively. Additionally, when $\alpha=0$, the solution represent the horizon of RN BH. Conversely, setting $q=0$ simplifies the solution of the Schwarzschild BH. Notably, in EMS model, the horizons emerge strictly under the conditions $0<q<\sqrt{2}$ and $0<\beta<\frac{(2-q^{2})^{2}}{4q^{2}}$ 10 ; 45 .
Figure 1: The above three pictures plot the function $f(r)$ for different values of $\beta$,$\alpha$ and $q$.
From the above two pictures, we can see there are two intersection points with $y=0$ which means the existence of the inner and outer horizons of BH. However, we see the third picture (EMD), when $\alpha>1$, there is only one intersection point with $y=0$. When the parameter $\alpha$ lies within the interval $(0,1)$, the EMD black hole possesses two horizons. Furthermore, we find that for a fixed $q$, an increase in $\alpha$ causes the function’s minimum value to decrease, whereas an increase in $\beta$ raises the minimum value. This may be attributed to the specific structure of the metric.
III Null geodesics and the Shadow
In this section, we will first explore the behavior of the shadow radius theoretically, and then find the constraints to the EMS parameter using the data from EHT. To begin with, consider a static, spherically symmetric spacetime given by
$ds^{2}$ $=g_{\mu\nu}dx^{\mu}dx^{\nu}$ (7)
$=-A(r)dt^{2}+B(r)dr^{2}+C(r)d\Omega^{2}$
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ is the line element of the unit two-spheres. Without loss of generality, we analyze the null geodesic in the equatorial plane only such that the polar angle is fixed to $\theta=\pi/2$. Then, the Hamiltonian for light ray is given by
$H=\frac{1}{2}g^{ik}p_{i}p_{k}=\frac{1}{2}\left(-\frac{p_{t}^{2}}{A(r)}+\frac{p_{r}^{2}}{B(r)}+\frac{p_{\phi}^{2}}{C(r)}\right).$ (8)
It is worth noting that since we consider EMS BH, the metric functions satisfy $A(r)=f(r)$, $B(r)=f(r)^{-1}$, and $C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}}$. The equations of motion for null particles are
$\left\{\begin{split}&\dot{x}^{i}=\frac{\partial H}{\partial p_{i}},\\ &\dot{p}_{i}=-\frac{\partial H}{\partial x^{i}}.\\ \end{split}\right.$ (9)
Here, $\dot{x}=dx/d\lambda$ and $\dot{p}$ represents the conjugate momenta. Eq.(9) gives
$\left\{\begin{split}&\dot{t}=-\frac{p_{t}}{f(r)},\quad\dot{\phi}=\frac{p_{\phi}}{C(r)},\quad\dot{r}=-\frac{p_{r}}{f(r)^{-1}},\\ &\dot{p}_{t}=0,\quad\dot{p}_{\phi}=0,\\ &\dot{p}_{r}=\frac{1}{2}\left(-\frac{p_{t}^{2}f^{\prime}(r)}{f(r)^{2}}+\frac{p_{r}^{2}f^{\prime}(r)^{-1}}{f(r)^{-2}}-\tilde{C}p_{\phi}^{2}\right).\\ \end{split}\right.$ (10)
where $\tilde{C}\equiv\frac{d}{dr}\frac{1}{C(r)}=-\frac{2\left(1-\frac{b_{2}}{r}\right)^{-\frac{2\alpha^{2}}{1+\alpha^{2}}}}{r^{3}}-\frac{2\alpha^{2}b_{2}\left(1-\frac{b_{2}}{r}\right)^{-1-\frac{2\alpha^{2}}{1+\alpha^{2}}}}{\left(1+\alpha^{2}\right)r^{4}}$. Setting $H=0$, we have
$-\frac{p_{t}^{2}}{f(r)}+\frac{p_{r}^{2}}{f(r)^{-1}}+\frac{p_{\phi}^{2}}{C(r)}=0,$ (11)
and it now follows that
$\frac{dr}{d\phi}=\frac{\dot{r}}{\dot{\phi}}=\frac{C(r)}{f(r)^{-1}}\frac{p_{r}}{p_{\phi}}.$ (12)
Setting $p_{t}=-\omega_{0}$, and using $p_{r}$, we get the relation how $r$ changes with $\phi$:
$\frac{dr}{d\phi}=\pm\frac{r^{2}}{C(r)^{1/2}f(r)^{-1/2}}\sqrt{\frac{\omega^{2}_{0}}{p^{2}_{\phi}}h(r)^{2}-1},$ (13)
where $h(r)^{2}=\frac{C(r)}{f(r)}$ is defined.
For a circular light orbit, the radial velocity and acceleration should be $\dot{r}=0$ and $\ddot{r}=0$ respectively, and hence, $p_{r}=0$. Eq.(11) then becomes
$0=-\frac{w^{2}_{0}}{f(r)}+\frac{p^{2}_{\phi}}{C(r)}.$ (14)
Since $\dot{p}_{r}=0$, Eq.(14) can be rewritten as
$\dot{p}_{r}=0=\frac{\omega_{0}^{2}f^{\prime}(r)}{f(r)^{2}}+\tilde{C}p^{2}_{\phi}.$ (15)
Using Eq.(14) and (15), we find
$p^{2}_{\phi}=C(r)\frac{\omega^{2}_{0}}{f(r)},$ (16)
and
$p^{2}_{\phi}=-\frac{\omega_{0}^{2}f^{\prime}(r)}{\tilde{C}f(r)^{2}}.$ (17)
The implication of subtracting Eqs. (16) and (17) give the information on how to find the radius of the photon sphere 176 ; 184
$0=\frac{d}{dr}h(r)^{2}.$ (18)
Figure 2: The above four pictures plot the photon sphere radius under the different parameters $\alpha$, $\beta$ and $q$.
Following Eq.(18), the analytical form of the photon sphere radius $r_{\text{ph}}$ is so complicated. We plot the above equation numerically (See Fig.2), where it shows the values of $\alpha$, $\beta$ and $q$ for the photon sphere to exist. When $q=0.5$, as the parameter $\alpha$ increases, the photon sphere radius essentially increases. Then, for a given value of $\alpha$, we observe an decrease in photon sphere radius as the $\beta$ increases. The second picture is exactly opposite from the first picture. When $\alpha=0$, as the parameter $q$ increases, the photon sphere radius essentially decreases. As $q$ approaches 0, the photon sphere radius asymptotically tends to 3 (Schwarzschild BH). Then, for a given value of $q$, we can observe an increase in photon sphere radius as the $\beta$ decreases. And when $\beta=0$, as the parameter $q$ increases, the photon sphere radius essentially decreases. As $q$ approaches 0, the photon sphere radius asymptotically tends to 3. Then, for a given value of $q$, we can observe an increase in photon sphere radius as the $\alpha$ increases.
Let us now determine the behavior of the shadow radius. For constructing the shadow, we assume that a static observer at radius coordinate $r_{\text{o}}$ sends light rays into the past. As can be seen from Fig.3, the angle $a$ between such a light ray and the radial direction is given by
$\cot a=\left.\frac{\sqrt{g_{rr}}}{\sqrt{g_{\phi\phi}}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}=\left.\frac{\sqrt{B(r)}}{\sqrt{C(r)}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}$ (19)
With the help of Eq.(12), we obtain
$\cot^{2}a=\frac{h\left(r_{\mathrm{O}}\right)^{2}}{h(R)^{2}}-1.$ (20)
By elementary trigonometry, we get 176
$\sin^{2}a=\frac{h(R)^{2}}{h\left(r_{\mathrm{O}}\right)^{2}},$ (21)
where $R$ means the minimum radius that the light ray approaches the center and then goes out.
The boundary curve of the shadow corresponds to past-oriented light rays that asymptotically approach one of the unstable circular light orbits at radius $r_{\mathrm{ph}}$. Therefore we have to consider the limit $R\rightarrow r_{\mathrm{ph}}$ in Eq.(21) for getting the angular radius $a_{\text{sh }}$ of the shadow 184 ,
$\sin^{2}a_{\mathrm{sh}}=\frac{h\left(r_{\mathrm{ph}}\right)^{2}}{h\left(r_{\mathrm{O}}\right)^{2}}.$ (22)
Here $h(r)$ is well-defined, $h(r)^{2}=C(r)/A(r)$. Note that the critical value $b_{\text{c}}$ of the impact parameter is connected with $r_{\mathrm{ph}}$ by Ref.176
$b_{\mathrm{c}}=h\left(r_{\mathrm{ph}}\right).$ (23)
Therefore we can also write Eq.(22) as
$\sin^{2}a_{\mathrm{sh}}=\frac{b_{\mathrm{c}}^{2}}{h\left(r_{\mathrm{O}}\right)^{2}},\quad\text{ or }\quad\sin^{2}a_{\mathrm{sh}}=\frac{b_{\mathrm{c}}^{2}A\left(r_{\mathrm{O}}\right)}{C\left(r_{\mathrm{O}}\right)}.$ (24)
For a general static spherically symmetric asymptotically flat spacetime, Ref.176 demonstrates that when the observer’s radial coordinate $r_{o}$ is sufficiently large, the black hole shadow radius approaches $b_{c}$. This radius is determined by the location of the photon sphere and the specific form of the metric function $f(r)$. In our paper, we will obey this consideration.
Figure 3: Example of calculation of light ray emitted from the observer’s position (small disk) into the past under an angle $a$. The BH horizon and the photon sphere are shown; $r_{\mathrm{ph}}$ is the photon sphere radius. The trajectory is calculated in the Asymptotically flat spacetime, $R$ denotes the radius coordinate at the point of closest approach. The picture is indicated in Ref.176 .
III.1 Constraints on parameters with the EHT observations of M87* and Sgr A*
In this subsection, we find constraints to the coupling parameters $\alpha$ and $\beta$ using the observation data provided by the EHT for M87* and Sgr A*. We only focused on the non-rotating case since the rotation parameter of Sgr A* is small enough to have considerable deviation to the shadow radius 110 . Furthermore, it has also been concluded for M87* that it is difficult to distinguish between Kerr and dilaton BHs based on BH shadow images alone using general-relativistic magnetohydrodynamical simulations and radiative-transfer calculations to generate synthetic shadow images in comparison to the present observation from the EHT 111 . Also remarked in Ref.112 that the shadow size of M87* lies within the range of $3\sqrt{3}(1\pm 0.17)M$, whether their model is spherically symmetric or axisymmetric.
[…]
$d_{\mathrm{sh}}=\frac{D\theta}{M}$ (25)
Figure 4: These plots are showing the constraints for different coupling parameters $\alpha$ and $\beta$.
Hence, the theoretical shadow diameter, however, can be obtained via
$d_{\text{sh }}^{\text{theo }}=2R_{\text{sh}}.$ (26)
Therefore, by using the above expression, we get the diameter of the shadow of M87* and Sgr A*
$\left\{\begin{split}&d_{\mathrm{sh}}^{\mathrm{M}87^{*}}=(11\pm 1.5)M,\\ &d_{\mathrm{sh}}^{\mathrm{Sgr}.\mathrm{A}^{*}}=(9.5\pm 1.4)M.\\ \end{split}\right.$ (27)
M87*$/\alpha$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$ M87*/$\beta$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$
charge $Q$ upper lower upper lower charge $Q$ upper lower upper lower
0.1 M - - - - 0.1M 99.00 - 45.32 -
0.2 M - - - - 0.2M 24.00 - 10.58 -
0.3 M - - - - 0.3M 10.11 - 4.149 -
0.4 M - - - - 0.4M 5.255 - 1.899 -
0.5 M - - - - 0.5M 3.008 - 0.8592 -
0.6 M - - - 3.441 0.6M 1.790 - 0.2960 -
0.7 M - - - 9.889 0.7M 1.058 - - -
0.8 M - - - 22.59 0.8M 0.5866 - - -
TABLE I. The table lists the values of $\alpha$ and $\beta$ at the $1\sigma$ and $2\sigma$ confidence levels for M87*.
Sgr A*$/\alpha$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$ Sgr A*/$\beta$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$
charge $Q$ upper lower upper lower charge $Q$ upper lower upper lower
0.1 M - - - - 0.1M - - 96.42 -
0.2 M - - - - 0.2M - - 23.36 -
0.3 M - - - - 0.3M - - 9.827 -
0.4 M - - - - 0.4M - - 5.094 -
0.5 M - - - - 0.5M - - 2.905 -
0.6 M - - - - 0.6M - - 1.718 -
0.7 M - - - - 0.7M - - 1.005 -
0.8 M - - - - 0.8M - - 0.5457 -
TABLE II. The table lists the values of $\alpha$ and $\beta$ at the $1\sigma$ and $2\sigma$ confidence levels for Sgr A*.
The variation of the diameter of the shadow image with coupling parameter $\alpha$ and $\beta$ for M87* and for Sgr A* is shown in Fig.4 , showing uncertainties at $1\sigma$ and $2\sigma$ levels. The numerical values for the upper or lower bounds in $\beta$ and $\alpha$ is found in Table I and Table II. However, it turns out that the data for M87* gives a better constraint for the coupling parameter $\beta$. It would mean that there is a certain value for the coupling parameter that gives the observed value of M87* shadow. For the case of Sgr A*, we find that the parameter $\alpha$ remains unconstrained within the $1\sigma$ and $2\sigma$ confidence intervals, whereas the parameter $\beta$ is well constrained within the $1\sigma$ interval.
III.2 Light bending: direct emission, lensing ring and photon ring
To gain a more comprehensive understanding of the appearance of EMS BH, we analyze its photon rings, and lensing rings in the presence of an optically bright accretion disk 177 ; 178 ; 179 . We begin by investigating the trajectory of light rays propagating in the vicinity of the EMS BH. To facilitate this analysis, we introduce the transformation $u=1/r$, which reformulates the orbit equation as follows
$(\frac{du}{d\phi})^{2}=G(u),$ (28)
where
$\sqrt{G(u)}=\frac{1}{C(1/u)^{1/2}f(1/u)^{-1/2}u^{2}}\sqrt{\frac{\omega^{2}_{0}}{p^{2}_{\phi}}h(1/u)^{2}-1}.$ (29)
For a static and spherically symmetric spacetime, we consider the motion of a particle confined to the equatorial plane. From the Lagrangian, two conserved quantities can be obtained from Eq. (10).
The impact parameter is defined as $b\equiv\frac{p_{t}}{p_{\phi}}$. When $b>b_{c}$, a light ray originating from infinity approaches the black hole, reaches a closest approach, and then escapes back to infinity. For $b<b_{c}$, the light ray inevitably falls into the black hole. In the critical case $b=b_{c}$, the light ray asymptotically orbits the black hole at the photon sphere radius $r_{\mathrm{ph}}$.
For $b>b_{c}$, the turning point of the trajectory corresponds to the smallest positive real root of $G(u)=0$, denoted as $u_{m}$. According to Eq. (28), the total change in the azimuthal angle $\phi$ along a trajectory with a given impact parameter $b$ can be expressed as
$\phi=2\int_{0}^{u_{m}}\frac{du}{\sqrt{G(u)}},\quad b>b_{c}.$ (30)
For $b<b_{c}$, we only focus on the trajectory outside the horizon, so the total change of azimuthal angle $\phi$ is obtained by
$\phi=\int_{0}^{u_{0}}\frac{du}{\sqrt{G(u)}},\quad b<b_{c}$ (31)
where $u_{0}=1/r_{+}$.
Figure 5: The behavior of photon trajectories around the Schwarzschild, RN, and EMS black holes as a function of the impact parameter b. In the upper panel, we present the total number of orbits, defined as $n=\phi/2\pi$. The trajectories are categorized based on $n$, where direct emission $n<3/4$ is shown in black, lensed trajectories 3/$43/4<n<5/$ 4 in yellow, and photon ring trajectories $n>5/4$ in red. The lower panel displays selected photon trajectories in Euclidean polar coordinates $(r,\phi)$. The spacing in the impact parameter is set to $1/10$, $1/100$, and $1/1000$ for the direct, lensed, and photon ring trajectories, respectively. The BH is represented as a solid disk, while the dashed black circle in the ray-tracing diagram marks the photon orbit. For the three cases studied, we set $\alpha=0,\beta=0,q=0$ for the Schwarzschild BH (first column), $\alpha=0,\beta=0,q=0.5$ for the RN BH (second column), and $\alpha=0.5,\beta=0.8,q=0.5$ for the EMS BH (third column).
To analyze the observational signatures of radiation emitted from the vicinity of a BH accretion disk. The Ref.117 divide trajectories into direct, lensed and photon rings. One define the total number of orbits $n=\frac{\phi}{2\pi}$ which is obviously a function of impact parameter $b$. We denote the solution of
$n(b)=\frac{2m-1}{4},\quad m=1,2,3,\cdots$ (32)
by $b_{m}^{\pm}$. Note that $b_{m}^{-}<b_{c}$ and $b_{m}^{+}>b_{c}$. Then we can classify all trajectories as follows
  • direct: $\frac{1}{4}<n<\frac{3}{4}\Rightarrow b\in\left(b_{1}^{-},b_{2}^{-}\right)\cup\left(b_{2}^{+},\infty\right)$
  • lensed: $\frac{3}{4}<n<\frac{5}{4}\Rightarrow b\in\left(b_{2}^{-},b_{3}^{-}\right)\cup\left(b_{3}^{+},b_{2}^{+}\right)$
  • photon ring: $n>\frac{5}{4}\Rightarrow b\in\left(b_{3}^{-},b_{3}^{+}\right)$
The physical picture of this classification is clear from the trajectory plots in Fig.5. Assuming light rays emit from north pole direction (far right of the trajectory plots), trajectories whose number of orbits 1/$41/4<n<3/$ 4 will intersect the equatorial plane only once. Trajectories whose number of orbits 3/$43/4<n<5/$ 4 will intersect the equatorial plane twice. Trajectories whose number of orbits $n>5/4$ will intersect the equatorial plane at least 3 times.
Parameter $\alpha=0,\beta=0,q=0$ $\alpha=0,\beta=0,q=0.5$ $\alpha=0.5,\beta=0.8,q=0.5$
Direct Emission $b<5.01514$ $b<4.77294$ $b<4.55452$
$n<3/4$ $b>6.16757$ $b>5.97448$ $b>5.81006$
Lensing Ring $5.187815.01514<b<$ 5.18781 $4.957934.77294<b<$ 4.95793 $4.753924.55452<b<$ 4.75392
3/$43/4<n<5/$ 4 $6.167575.22794<b<$ 6.16757 $5.974485.0039<b<$ 5.97448 $5.810064.80685<b<$ 5.81006
Photon Ring
$n>5/4$ $5.227945.18781<b<$ 5.22794 $5.00394.95793<b<$ 5.0039 $4.806854.75392<b<$ 4.80685
TABLE III. The region of direct emission, lensing ring, and photon ring for the different parameter $\alpha,\beta,q$.
Table III illustrates the variation of the BH shadow with increasing parameters $\alpha$, $\beta$, and $q$. From the table and corresponding figures, it is observed that the range of the lensing and photon rings expands progressively from the Schwarzschild BH to the RN BH and further to the EMS BH. This indicates that, from left to right, the contribution to the brightness of the lensing and photon rings increases accordingly. Moreover, when the impact parameter approaches the critical value $b\to b_{c}$, the photon orbit exhibits a sharp peak in the $(b,\phi)$ plane. Beyond this regime, as $b$ further increases, the photon trajectories correspond predominantly to direct emission across all cases considered.
III.3 Transfer functions and observed specific intensities
We analyze the emitted intensity from the EMS BH, assuming that the disk radiates isotropically in the rest frame of a static observer. According to Liouville’s theorem, the quantity $I_{\nu}^{\mathrm{em}}/\nu_{e}^{3}$ remains conserved along the trajectory of the light ray. Consequently, the observed intensity can be expressed as
$I_{\nu^{\prime}}^{\mathrm{obs}}=g^{3}I_{\nu}^{\mathrm{em}},$ (33)
[…]
$I^{\mathrm{obs}}=g^{4}I^{\mathrm{em}}.$ (34)
$I_{\mathrm{em}}$ here is the total emitted specific intensity from the accretion disk. Therefore, the total intensity received by the observer will be
$I^{\mathrm{obs}}(b)=\left.\sum_{n}g^{4}I^{\mathrm{em}}(r)\right|_{r=r_{m}(b)},$ (35)
where $r_{m}(b)$ is the $m$ intersection outside the horizon in the equatorial plane, which we call the transfer function. The transfer function establishes the connection between the radial coordinates and the photon’s impact parameter. It is important to note that our analysis does not account for light absorption or reflection by the accretion disk, nor does it consider the attenuation of light intensity due to environmental effects. Instead, we work within an idealized model.
We denote the solution of the orbit equation by $u(\phi,b)$ and focus on the first three transfer functions, which can be obtained by
$\left\{\begin{split}&r_{1}(b)=\frac{1}{u\left(\frac{\pi}{2},b\right)},&b\in\left(b_{1}^{-},\infty\right)\\ &r_{2}(b)=\frac{1}{u\left(\frac{3\pi}{2},b\right)},&b\in\left(b_{2}^{-},b_{2}^{+}\right)\\ &r_{3}(b)=\frac{1}{u\left(\frac{5\pi}{2},b\right)},&b\in\left(b_{3}^{-},b_{3}^{+}\right)\end{split}\right.$ (36)
As described in Ref.120 , the first transfer function corresponds to the “direct image” of the disk, which primarily reflects the redshift of the source profile. The second transfer function produces a highly demagnified image of the disk’s far side, known as the “lensing ring”. The third transfer function yields an extremely demagnified image of the near side of the disk, referred to as the “photon ring”. Images generated by subsequent transfer functions are so severely demagnified that they can be effectively disregarded.
Figure 6: The first three transfer functions in BHs for different values of $\alpha,\beta,q$. The above figure, from left to right, respectively Schwarzschild BH, RN BH and EMS BH. Here, the y-axis is the transfer function $r_{m}(b)$ and the x-axis is the impact parameter b.
In Fig.6, the black dots $(m=1)$ represent the transfer function for direct emission and exhibit an almost constant slope, indicating a redshifted source profile. The yellow dots $(m=2)$ correspond to the lensing ring, where the slope remains small near $b_{c}$ but increases rapidly with $b$, leading to significant demagnification of the image from the far side of the accretion disk. The red dots $(m=3)$ represent the photon ring, where the slope approaches infinity, implying extreme demagnification of the disk’s near-side image. Consequently, the dominant contribution to the observed flux arises from direct emission, while higher-order images $(m>3)$ contribute negligibly and can be disregarded.
All three black holes exhibit these characteristics. The critical impact parameters are $b_{c}=5.19615$ for the Schwarzschild BH, $b_{c}=4.96791$ for the RN BH, and $b_{c}=4.76598$ for the EMS BH. This result indicates that the Schwarzschild BH has the largest critical impact parameter, corresponding to the largest shadow radius, while an increase in the model parameters leads to a reduction in the BH’s shadow size.
III.4 Observational features of direct emission, photon and lensing rings
As discussed in the previous section, for an observer at infinity, the specific intensity depends solely on the radial coordinate $r$. In our analysis, we consider three toy models for the emissivity profile $I^{\mathrm{em}}$.
With the transfer function determined, we now consider a specific emission profile. As the first model, we examine a scenario in which the emission is sharply concentrated at the innermost stable circular orbit (ISCO) given by
$I_{1}^{\mathrm{em}}(r)=\begin{cases}\left(\frac{1}{r-\left(r_{\mathrm{isco}}-1\right)}\right)^{2},&r>r_{\mathrm{isco}}\\ 0,&r\leq r_{\mathrm{isco}}\end{cases}$ (37)
Second model, we consider the emission is sharply peaked at the photon sphere, and it ends abruptly at $r=r_{\text{ph}}$ while quickly decaying to zero, such as
$I_{2}^{\mathrm{em}}(r)=\begin{cases}\left(\frac{1}{r-\left(r_{\mathrm{ph}}-1\right)}\right)^{3},&r>r_{\mathrm{ph}}\\ 0,&r\leq r_{\mathrm{ph}}\end{cases}$ (38)
Final model, we consider an emission decaying gradually from the horizon to the ISCO, such as
$I_{3}^{\mathrm{em}}(r)=\begin{cases}\frac{1-\arctan\left(r-\left(r_{\text{isco}}-1\right)\right)}{1-\arctan\left(r_{\mathrm{ph}}\right)},&r>r_{\mathrm{+}}\\ 0,&r\leq r_{\mathrm{+}}\end{cases}$ (39)
Each of these three models exhibits distinct characteristics. In particular, the second model features a rapid decay rate, whereas the third model exhibits a much slower decay. In the third model, the emission originates directly from the event horizon, while in the second model, it begins at the photon sphere, and in the first model, it is confined to the ISCO. Although these models represent highly idealized scenarios, they provide valuable qualitative insights into the behavior of photons around the BH.
Figure 7: The observational appearance of the thin disk with different emission profiles for $\alpha=0,\beta=0,q=0$ is shown from a face-on perspective. The first row corresponds to the emission profile intensity as described by model 1, the second row represents model 2, and the third row corresponds to model 3, as outlined in Section III.C. In the plots, the emitted and observed intensities, $I^{\mathrm{em}}$ and $I^{\mathrm{obs}}$, are normalized to the maximum value $I^{0}$ of the emitted intensity outside the event horizon.
Figure 8: The observational appearance of the thin disk with different emission profiles for $\alpha=0,\beta=0,q=0.5$ is shown from a face-on perspective. This figure presents the results of the three models corresponding to RN BH.
Figure 9: The observational appearance of the thin disk with different emission profiles for $\alpha=0.5,\beta=0.8,q=0.5$ is shown from a face-on perspective. This figure presents the results of the three models corresponding to EMS BH.
In the Fig.7, 8, and 9, we have shown the observed appearance of the thin accretion disk with the observed intensity and the impact factor corresponding to the three models for different coupling constant parameters respectively. In all the figures, we see that for the first model (top row), the emission intensity (top-first) has a peak near the critical impact parameter $r\sim b_{c}$ and then it decreases as radial distance increases and becomes zero. In this case, the photon sphere lies in the interior region of the emission part of the disk. As can be readily seen from the middle panel (top-second), the direct image of the disk looks very similar to the emission profile, although its abrupt end occurs at different place due to gravitational lensing. However, we noticed that the peak of the photon and lensing rings are not only smaller than the direct emission but also have a narrow observational area. Hence, we observed that the observed intensity has a huge contribution coming from the direct emission, a small contribution from the lensing rings, and very little contribution from the photon rings. As is evident from the right panel (2D shadow image, top-third) where the lensing ring emission appears as a tiny ring inside the direct image and one can barely see the photon ring emission in the right panel if one zooms in.
For the second model (middle row), the intensity of the emission has a peak at photon sphere $r\sim r_{\mathrm{ph}}$ and then decreases as radial distance increases (middle-first). In this case, the observed intensity profile (middle-second) peaks due to the direct emission, and then it shows an attenuation with increasing $r$. We observed that the photon and lensing rings, which improves the total intensity of this particular area, and therefore, we get a new peak due to the photon ring, lensing ring, and direct emission. However, the photon and lensing rings are highly demagnetized and have a very narrow area in the observed intensity. Therefore, we still have a dominant contribution coming from the direct emission in the observed intensity which can be seen in the $2\mathrm{D}$ shadow image (middle-third).
For the third model (bottom row), the peak in the intensity starts from the horizon ($r_{+}$) and then decreases with increasing $r$ (bottom-first). In this case, redshift effects noticeably decrease the observed flux. However, the most important difference from the top two row for our considerations is that the lensing ring and photon ring emission are now superimposed on the direct emission. However, as discussed in the last section, the photon and lensing rings are highly demagnetized and have a very narrow area in the observed intensity. Therefore, we still have a dominant contribution coming from the direct emission in the observed intensity which can be seen in the 2D shadow image (bottom-third).
Although Fig.7, 8 and 9 show only a few highly idealized cases of thin disk emission near a BH (viewed face on), it illustrates two key points that we believe will hold quite generally for optically thin disk emission: (1) The emission is dominated by the direct emission, with the lensing ring emission providing only a small contribution to the total flux and the photon ring providing a negligible contribution in all cases. (2) For all three models, as we increase the coupling constant $\alpha,\beta$ and $q$, the observed intensity decreases, and it is much less than what we get for the Schwarzschild BH.
IV Conclusions
In this paper, we investigated the observational characteristics of thin accretion disk surrounding the EMS BH. We first analyzed the conditions under which the BH possesses inner and outer horizons and derived the theoretical parameter range. Furthermore, we examined the geodesic equation of photons and determined the location of the photon sphere for the EMS BH. Our results indicate that as the parameter $\alpha$ increases, the radius of the photon sphere enlarges, whereas an increase in $\beta$ leads to a decrease in the photon sphere radius. Subsequently, employing the analytical expression for the shadow of spherically symmetric BH, we constrained the model parameters $\alpha$ and $\beta$ using the angular diameter distance from the EHT data within the $1\sigma$ and $2\sigma$ confidence intervals. Our findings reveal that the parameter $\beta$ is well constrained, while the constraints on $\alpha$ remain less stringent.
On the other hand, we analyzed the photon trajectories of the Schwarzschild BH, RN BH, and EMS BH. The differences among these three types of BHs are only reflected in the critical impact parameter $b_{c}$ (the BH shadow radius). We found that the Schwarzschild BH has the largest shadow radius, while the EMS BH has the smallest. All three types of BHs undergo a transition from direct emission to the photon ring, then to the lensing ring, and finally back to direct emission as the impact parameter $b$ increases. A similar effect has been observed using the Okyayvg Mathematical Notebook package 182 in our investigation of BH shadows and rings with three toy models of thin accretion disk. Our analysis reveals that the observed intensity is primarily dominated by direct emission, while the lensing ring contributes only marginally to the total flux, and the photon ring remains negligible in all cases. Among the three types of BHs, the Schwarzschild BH exhibits the highest emission intensity, followed by the RN BH, with the EMS BH having the lowest intensity. In other words, an increase in the model parameters leads to a reduction in the disk’s radiation intensity.
Acknowledgements.
This work was supported by the National Key Research and Development Program of China (No. 2022YFA1403700), NSFC (Grants No. 12334002), Guangdong Provincial Quantum Science Strategic Initiative Grand No. SZZX2401001, the Science, Technology and Innovation Commission of Shenzhen Municipality (No. ZDSYS20190902092905285), and Center for Computational Science and Engineering at Southern University of Science and Technology.

### v3

@@
 These observations offer valuable constraints on the geometry of spacetime in the strong-field regime, particularly near the event horizon, allowing precise measurements of BH mass and spin 135 ; 136 ; 145 ; 146 ; 147 ; 148 ; 149 ; 150 ; 151 ; 152 ; 153 ; 154 ; 155 ; 185 .
-Collectively, these findings provide a crucial platform for testing GR and exploring potential modifications to gravitational theories in extreme astrophysical environments 137 .
+Collectively, these findings provide a crucial platform for testing GR and exploring potential modifications to gravitational theories in extreme astrophysical environments 137 ; Liu:2024brf ; Yin:2025coq ; He:2022lrc ; Wu:2024sng ; Feng:2024iqj ; Feng:2023iha ; Feng:2022bst .
 It is well known that GR has demonstrated remarkable accuracy in various tests 138 .
@@
 Among various modified gravity models, the Einstein-Maxwell-Scalar (EMS) model serves as a promising candidate.
-Since the EMS theory allows for hairy BH solutions 143 ; 144 ; 145 , where BHs can possess scalar hair, investigating their properties in the strong-field regime is of significant interest.
+In fact, various specific realizations of EMS models naturally arise as consistent truncations or low-energy limits of string theory.
+For instance, in heterotic string theory, compactification from ten to four dimensions in the Einstein frame yields an effective action containing the metric, a dilaton-like scalar field, and Abelian gauge fields with an exponential coupling of the form $f(\phi)F_{\mu\nu}F^{\mu\nu}$ — as in the well-known Gibbons–Maeda–Garfinkle–Horowitz–Strominger (GMGHS) solution Gibbons:1988 ; Garfinkle:1990qj .
+Similar structures occur in Type II string compactifications, where moduli fields (Kähler, complex structure, axion–dilaton) couple to gauge fields descending from the NS–NS and R–R sectors Polchinski:1998 ; Becker:2006 .
+In extended supergravity, particularly $\mathcal{N}=4$ and $\mathcal{N}=8$ theories, the gauge kinetic matrix $f_{IJ}(\phi)$ depends explicitly on scalar fields Cremmer:1979 ; deWit:1982 .
+Truncation to a single $U(1)$ vector and one scalar degree of freedom yields precisely the EMS form.
+The $\mathcal{N}=2$ STU model Duff:1996 provides another example, in which selecting a single scalar–vector sector produces an EMS-type action.
+The EMS action serves as a powerful unifying framework, encompassing two distinct and mutually exclusive branches of black hole solutions: the Einstein–Maxwell–Dilaton (EMD) branch and the spontaneously scalarized Reissner–Nordström (RN) branch.
+The profound differences between these two scenarios are encoded solely in the local behavior of the coupling function $K(\phi)$ at $\phi=0$.
+The EMD limit, realized by choosing an exponential coupling $K(\phi)=\mathrm{e}^{2\alpha\phi}$, is characterized by $K(0)=1$ and $K^{\prime}(0)\neq 0$.
+This condition renders the electro-vacuum RN solution, where $\phi\equiv 0$, an invalid background.
+Consequently, EMD black holes are born with an intrinsic, non-trivial scalar field, often referred to as “dilaton hair” and they never bifurcate from a hairless RN seed.
+Their asymptotic behavior is distinct, approaching a dilaton asymptotic at infinity, and they can violate the RN extremal bound on the charge-to-mass ratio Garfinkle:1990qj ; belkhadria_2025 .
+In contrast, the spontaneously scalarized RN branch emerges from a different class of coupling functions satisfying $K(0)=1$, $K^{\prime}(0)=0$, and $K^{\prime\prime}(0)<0$.
+This specific set of conditions permits the RN metric with $\phi\equiv 0$ as a legitimate and stable background in the absence of a strong electric field.
+However, in the presence of a sufficiently strong electric charge, a tachyonic instability can arise.
+The effective mass squared of the scalar field, $\mu_{\text{eff}}^{2}(r)=K^{\prime\prime}(0)\,Q^{2}/r^{4}$, becomes negative within a certain radial window.
+This instability triggers a second-order phase transition, where the hairless RN black hole spontaneously acquires scalar hair, bifurcating from the original RN branch at a critical charge Qiu_2021 .
+Asymptotically, these scalarized black holes return to the RN solution, as the scalar hair vanishes at infinity.
+Since the EMS theory allows for hairy BH solutions 143 ; 144 ; 145 ; 189 , where BHs can possess scalar hair, investigating their properties in the strong-field regime is of significant interest.
 Previous studies have shown that dilaton BH spacetimes are influenced by the dilaton charge, which not only serves as a new form of hair but also provides a powerful tool for analyzing optical phenomena in such backgrounds.
-In this work, we consider a charged BH solution within the EMS theory as an extension of the Reissner-Nordström (RN) solution.
+In this work, we consider a charged BH solution within the EMS theory as an extension of the RN solution.
 Unlike dilaton BHs, where the scalar field typically contributes through an effective charge, the BH solution in EMS theory is directly governed by its model parameters, which serve as the primary distinguishing feature.
@@
 Furthermore, we investigated the observable effects of radiation emitted by the accretion disk and found that the EMS BH exhibits a significantly reduced radiation intensity compared to the Schwarzschild and RN BHs.
-Finally, we present our conclusions and provide an outlook in Sec.
-IV.
-The paper has been written with the unit system $G=M=c=1$, and metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ).
+And in Sec III.E, we also explore the deflection angle in the weak gravitational lensing.
+Finally, we present our conclusions and provide an outlook in Sec.IV.
+The paper has been written with the unit system $G=M=c=1$, and metric signature (,+,+,+)(-,+,+,+).
 II The black hole solution in Einstein-Maxwell-scalar theory
-The interaction between a scalar field and gravity was first systematically studied by Fisher, who obtained a static and spherically symmetric solution to the Einstein equations in the presence of a massless scalar field 102 .
-This pioneering work laid the foundation for subsequent investigations into the role of scalar fields in gravitational theories.
-In particular, the EMS theory has garnered considerable interest due to its profound physical implications and its natural emergence in various theoretical frameworks.
-Notably, it arises in the context of Kaluza-Klein theories, where extra-dimensional reductions lead to scalar field interactions 103 .
-Furthermore, it plays a crucial role in supergravity and string theory, where dilaton fields naturally couple to gravity and gauge fields 104 .
-Additionally, scalar fields are fundamental in cosmology, influencing early universe dynamics, inflationary models, and the late-time acceleration of the universe 105 .
-These diverse connections underscore the significance of the EMS theory in modern theoretical physics.
-Within the framework of effective field theory (EFT), the interaction between a scalar field and an electromagnetic field can be incorporated through specific coupling terms in the Lagrangian.
-Such couplings naturally arise in the low-energy effective description of string theory, where the scalar field is identified with the dilaton—a massless excitation resulting from the compactification of extra spatial dimensions—while the electromagnetic field originates from gauge fields in the higher-dimensional theory.
-The dilaton, as a key modulus field, encodes information about the compactification geometry and the associated moduli space, dynamically influencing the strength of gauge interactions in the four-dimensional effective theory.
-The coupling between the scalar and electromagnetic fields is an inevitable consequence of dimensional reduction, where the dilaton acts as a dynamical field that rescales the effective gauge couplings.
-This interaction not only modifies the propagation and self-interactions of electromagnetic waves but also introduces additional curvature terms in the gravitational sector, leading to non-trivial corrections to the Einstein field equations when the scalar field is present.
-These effects are particularly relevant in scenarios involving strong-field gravity, early-universe cosmology, and potential deviations from classical electrodynamics in high-energy regimes.
-Further implications include modifications to BH solutions, cosmological inflation models, and the possible emergence of new observational signatures in astrophysical and laboratory settings 169 ; 170 ; 171 ; 172 ; 173 ; 174 ; 175 ; 186 .
-Such interaction terms are often introduced to match experimental or astrophysical observations under specific conditions. A typical form of the interaction term is given by
-$\mathcal{L}_{\text{int}}=f(\phi)F_{\mu\nu}F^{\mu\nu},$ (1)
-where $f(\phi)$ is a function of the scalar field $\phi$, and $F_{\mu\nu}$ represents the electromagnetic field strength tensor.
-Depending on the functional form of $f(\phi)$, the coupling may be constant, exponential, or take other more complex forms.
-These interaction terms can significantly influence the propagation of electromagnetic fields, especially in regions with strong scalar field backgrounds, such as near BHs or on cosmological scales.
-Furthermore, the inclusion of a scalar field can be interpreted as a modification to standard Einstein gravity.
-Such frameworks are often categorized as scalar-tensor theories or extended theories of gravity.
-In these models, the coupling between the scalar field and the electromagnetic field can significantly influence the dynamics of the electromagnetic field and modify the solutions of the gravitational field equations.
-For example, investigations within the EMS theory have demonstrated that the presence of a scalar field can alter the electromagnetic properties of BHs, giving rise to novel classes of BH solutions, such as those characterized by scalar-electromagnetic coupling.
-We consider the EMS theory described by the following action
-$S[g_{\mu\nu},\phi]=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}(R-2\nabla_{\mu}\phi\nabla^{\mu}\phi-K(\phi)F_{\mu\nu}F^{\mu\nu}-V(\phi)),$ (2)
-where $R$ is the Ricci scalar.
-The EMS theory describes the real scalar field $\phi$ minimally coupled to GR and non-minimally coupled to Maxwell’s background.
-$K(\phi)$ is the coupling function between the Maxwell and scalar field, and $V(\phi)$ is the scalar potential.
-if we set up the determined expressions of $K(\phi)$ and $V(\phi)$ first, then the corresponding BH solutions are determined.
-For example, if we set $K=1$ and $V=2\lambda$, $\lambda$ is the cosmological constant, then the EMS theory gives the Reissner-Nordstrom-de-Sitter solution.
-And if we set $K=e^{2\phi}$ and $V=0$, then we can get the dilation BH solution 1 ; 2 .
+The systematic investigation of scalar-gravity coupling was introduced by Fisher, who obtained static spherically symmetric solutions to Einstein’s equations with a massless scalar field 102 .
+This foundational work established the theoretical basis for subsequent developments, particularly in Einstein-Maxwell-scalar (EMS) theory which has attracted significant attention due to its implications and natural emergence in: (i) Kaluza-Klein dimensional reduction 103 , (ii) dilaton couplings in string theory 104 , and (iii) cosmological inflation and late-time acceleration 105 .
+The EMS action takes the form
+$S=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}\left[R-2(\nabla\phi)^{2}-K(\phi)F^{2}-V(\phi)\right],$ (1)
+where $R$ is the Ricci scalar, $\phi$ is the scalar field.
+$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ represents the electro-magnetic field tensor.
+The field equations derived from this action yield
+$\begin{cases}\nabla_{\mu}[K(\phi)F^{\mu\nu}]=0,\\ \Box\phi=\frac{1}{4}\left[\partial_{\phi}K(\phi)F^{2}+\partial_{\phi}V(\phi)\right],\\ R_{\mu\nu}=2\partial_{\mu}\phi\partial_{\nu}\phi+\frac{1}{2}g_{\mu\nu}V(\phi)+2K(\phi)(F_{\mu\sigma}F^{\sigma}_{~\nu}-\frac{1}{4}g_{\mu\nu}F^{2}).\end{cases}$ (2)
+The coupling function $K(\phi)$ and potential $V(\phi)$ determine the theory’s physical content.
 Some important solutions with different $K(\phi)$ and $V(\phi)$ are given by 3 ; 4 ; 5 ; 6 ; 7 ; 8 .
-In our research, we consider $V=0$ and the coupling function $K(\phi)$ is
-$K(\phi)=\frac{(\alpha^{2}+1)e^{\frac{-2\phi}{\alpha}}}{(\alpha^{2}+1+\beta)e^{\frac{-2\phi(\alpha^{2}+1)}{\alpha}}+\beta\alpha^{2}},$ (3)
-with $\alpha$ and $\beta$ as characteristic parameters or coupling constants; it is observed that when $\beta\to\infty$, the influence of the electromagnetic field is negligible.
-Conversely, in the extremal case where $\beta=0$, it simplifies to the Einstein-Maxwell-Dilaton (EMD) coupling described by $K(\phi)=e^{2\alpha\phi}$.
-The equations of motion are obtained by varying the action with respect to the metric, dilaton, and Maxwell field, respectively, as follows
-$\left\{\begin{array}[]{l}0=\nabla_{\mu}\left[K(\phi)F^{\mu\nu}\right],\\ \Box\phi=\frac{1}{4}\frac{\partial K(\phi)}{\partial\phi}F^{2},\\ R_{\mu\nu}=2\partial_{\mu}\phi\partial_{\nu}\phi+2K(\phi)\left(F_{\mu\sigma}F^{~{}\sigma}_{\nu}-\frac{1}{4}g_{\mu\nu}F^{2}\right).\end{array}\right.$ (4)
-Due to the field equations mentioned above, Ref.9 provides a BH solution, which can be described as
-$\left\{\begin{split}&ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+C(r)(d\theta^{2}+\sin^{2}\theta d\phi^{2})\\ &f(r)=\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{\frac{1-\alpha^{2}}{1+\alpha^{2}}}+\frac{\beta Q^{2}}{C(r)}\\ &C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}},\end{split}\right.$ (5)
-where $b_{1}$ and $b_{2}$ are functions solely of $\alpha$, yielding
-$\left\{\begin{split}b_{1}&=\left(1+\sqrt{1-q^{2}(1-\alpha^{2})}\right)M,\\ b_{2}&=\frac{1+\alpha^{2}}{1-\alpha^{2}}\left[1-\sqrt{1-q^{2}(1-\alpha^{2})}\right]M.\end{split}\right.$ (6)
-with $q\equiv{Q}/{M}$ the charge-to-mass ratio, and $M$ the mass of the BH.
-The location of horizons is defined where $f(r_{\pm})=0$.
-Specifically, for $\beta=0$, these configurations align with the GMGHS solutions, as referenced in 11 , where $r_{\pm}$ are identified as $b_{1}$ and $b_{2}$, respectively.
-Additionally, when $\alpha=0$, the solution represent the horizon of RN BH.
-Conversely, setting $q=0$ simplifies the solution of the Schwarzschild BH.
-Notably, in EMS model, the horizons emerge strictly under the conditions $0<q<\sqrt{2}$ and $0<\beta<\frac{(2-q^{2})^{2}}{4q^{2}}$ 10 ; 45 .
-Figure 1: The above three pictures plot the function $f(r)$ for different values of $\beta$,$\alpha$ and $q$.
+Notable special solutions include
+  • RN-de Sitter BH $K(\phi)=1$, $V=2\Lambda$, $\phi=0$,
+  • Dilaton BH $K(\phi)=e^{2\phi}$, $V=0$ 1 ; 2 .
+Different choices of the coupling function $K(\phi)$ lead to qualitatively distinct physics.
+Polynomial forms (such as $K(\phi)=1+\alpha\phi^{2}$) preserve the RN branch but allow for spontaneous scalarization above a critical threshold, where the scalar field effectively screens the electromagnetic charge and produces multiple scalarized branches.
+Exponential couplings (such as $K(\phi)=e^{2\alpha\phi}$) correspond to dilatonic models in which RN solutions no longer exist and charged black holes necessarily carry scalar hair, with modified extremality and thermodynamics.
+Other bounded or interpolating forms regulate the growth of scalar hair or bridge between EMS-type and dilatonic behavior.
+Our investigation focuses on the parameterized coupling function Yu_2021
+$K(\phi)=\frac{(\alpha^{2}+1)e^{-2\phi/\alpha}}{(\alpha^{2}+1+\beta)e^{-2\phi(\alpha^{2}+1)/\alpha}+\beta\alpha^{2}},$ (3)
+which exhibits two physically significant limits.
+(i) $\beta\to\infty$: Decoupled electromagnetic sector.
+(ii) $\beta\to 0$: Einstein-Maxwell-Dilaton (EMD) theory with $K(\phi)=e^{2\alpha\phi}$.
+(iii) $\beta\to 0$: Einstein-Maxwell-Dilaton (EMD) theory with $K(\phi)=e^{2\alpha\phi}$.
+The physical meaning of $\alpha$ is that it measures the strength of the coupling between the dilaton scalar field and the Maxwell field.
+It first emerges in the EMD limit ($\beta=0$) and is inherited from low-energy string theory, where the dilaton couples universally to gauge fields.
+Parameter $\beta$ controls the relative contribution of the RN and dilaton branches in the metric.
+Regarding its origin, it is introduced to combine the RN and dilaton line elements into a single interpolating family, thereby extending the EMD class.
+As shown in Ref.9 , these admit BH solution with metric components
+$\begin{cases}ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+C(r)d\Omega^{2},\\ f(r)=\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{\frac{1-\alpha^{2}}{1+\alpha^{2}}}+\frac{\beta Q^{2}}{C(r)},\\ C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}},\end{cases}$ (4)
+where the parameters $b_{1}$ and $b_{2}$ are determined by
+$\begin{cases}b_{1}=\left(1+\sqrt{1-q^{2}(1-\alpha^{2})}\right)M,\\ b_{2}=\frac{1+\alpha^{2}}{1-\alpha^{2}}\left[1-\sqrt{1-q^{2}(1-\alpha^{2})}\right]M,\end{cases}$ (5)
+with $q\equiv Q/M$ representing the dimensionless charge-to-mass ratio.
+Moreover, since obtaining analytic expressions for the inner and outer horizons of the EMS black hole is extremely difficult, we consider two special cases, while the general EMS case is determined numerically.
+(I) $\alpha=1$: With this definition established, the following relations can be derived Yu_2021
+$\begin{cases}r_{+}=M\sqrt{2-q^{2}-\beta q^{2}+\sqrt{\left(2-q^{2}\right)^{2}-4q^{2}\beta}},\\ r_{-}=M\sqrt{2-q^{2}-\beta q^{2}-\sqrt{\left(2-q^{2}\right)^{2}-4q^{2}\beta}}.\end{cases}$ (6)
+For analytical convenience, we take $Q>0$ without imposing any loss of generality. The condition ensuring the existence of both the inner and outer horizons can be demonstrated to take the form
+$0<q<\sqrt{2},\quad\quad\beta<\frac{\left(2-q^{2}\right)^{2}}{4q^{2}}.$ (7)
+Moreover, the condition for the existence of a single horizon can be written as
+$0<q<\sqrt{2},\quad\quad\beta=\frac{(2-q^{2})^{2}}{4q^{2}},$ (8)
+under which the two horizons coincide, giving rise to an extremal black hole.
+This demonstrates that an extremal configuration can occur once the coupling constant $\beta$ meets the above criterion, irrespective of how small the charge is.
+(2) $\beta=0$ (the EMD solution): In this case, the inner and outer horizons are determined by $b_{1}$ and $b_{2}$.
+[…]
+The horizons are determined by the black hole’s mass $M$, charge $Q$, and the dilaton coupling constant $\alpha$.
+For two distinct real horizons to exist, the condition $q^{2}\leq 1+\alpha^{2}$ must hold.
+The equality $q^{2}=1+\alpha^{2}$ corresponds to the extremal case, where the two horizons merge into a single degenerate horizon.
+In contrast, the condition $q^{2}>1+\alpha^{2}$ eliminates the horizons and exposes a naked singularity.
+Therefore, when $\alpha\neq 1$ with arbitrary $\beta$, or when $\beta=0$ with arbitrary $\alpha$, a case-by-case analysis is necessary to ensure that naked singularities do not emerge in the black hole solutions.
+Notably, both situations correspond to special subclasses of the EMS framework.
+In the following analysis, we confine the parameters to these well-defined ranges to guarantee the existence of black hole solution.
+Besides these, the black-hole horizons for all other parameter choices are already plotted in Figure 1.
+In the figure below, we plot the variation of the metric components of the EMS solution with respect to the coordinate radius $r$.
+Figure 1: The above three pictures plot the function $f(r)$ for different values of $\beta$$\alpha$ and $q$.
 From the above two pictures, we can see there are two intersection points with $y=0$ which means the existence of the inner and outer horizons of BH.
-However, we see the third picture (EMD), when $\alpha>1$, there is only one intersection point with $y=0$.
-When the parameter $\alpha$ lies within the interval $(0,1)$, the EMD black hole possesses two horizons.
+However, in the third case (EMD), we observe that both a single intersection and two intersections can occur.
+All such parameter choices lie within the admissible ranges discussed above, thereby ensuring the absence of naked singularities.
 Furthermore, we find that for a fixed $q$, an increase in $\alpha$ causes the function’s minimum value to decrease, whereas an increase in $\beta$ raises the minimum value.
@@
 III Null geodesics and the Shadow
-In this section, we will first explore the behavior of the shadow radius theoretically, and then find the constraints to the EMS parameter using the data from EHT.
-To begin with, consider a static, spherically symmetric spacetime given by
-$ds^{2}$ $=g_{\mu\nu}dx^{\mu}dx^{\nu}$ (7)
-$=-A(r)dt^{2}+B(r)dr^{2}+C(r)d\Omega^{2}$
-where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ is the line element of the unit two-spheres.
-Without loss of generality, we analyze the null geodesic in the equatorial plane only such that the polar angle is fixed to $\theta=\pi/2$.
-Then, the Hamiltonian for light ray is given by
+In this section, we will constrain the parameter of $\alpha$ and $\beta$ by comparing the theoretically defined shadow radius with observational data from the EHT.
+We begin by considering a static, spherically symmetric spacetime as presented in Ref.177
+$ds^{2}=-A(r)dt^{2}+B(r)dr^{2}+C(r)\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (10)
+The functions $A(r)$, $B(r)$, and $C(r)$ satisfy the asymptotic flatness conditions, with $A(r)$ and $B(r)$ $\rightarrow 1$ at spatial infinity while $C(r)\rightarrow r^{2}$.
+For simplicity, we restrict our analysis to null geodesics in the equatorial plane by fixing the polar angle at $\theta=\pi/2$.
+Under this configuration, the photon’s Hamiltonian takes the form
 $H=\frac{1}{2}g^{ik}p_{i}p_{k}=\frac{1}{2}\left(-\frac{p_{t}^{2}}{A(r)}+\frac{p_{r}^{2}}{B(r)}+\frac{p_{\phi}^{2}}{C(r)}\right).$ (11)
 It is worth noting that since we consider EMS BH, the metric functions satisfy $A(r)=f(r)$, $B(r)=f(r)^{-1}$, and $C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}}$.
-The equations of motion for null particles are
-$\left\{\begin{split}&\dot{x}^{i}=\frac{\partial H}{\partial p_{i}},\\ &\dot{p}_{i}=-\frac{\partial H}{\partial x^{i}}.\\ \end{split}\right.$ (9)
-Here, $\dot{x}=dx/d\lambda$ and $\dot{p}$ represents the conjugate momenta. Eq.(9) gives
-$\left\{\begin{split}&\dot{t}=-\frac{p_{t}}{f(r)},\quad\dot{\phi}=\frac{p_{\phi}}{C(r)},\quad\dot{r}=-\frac{p_{r}}{f(r)^{-1}},\\ &\dot{p}_{t}=0,\quad\dot{p}_{\phi}=0,\\ &\dot{p}_{r}=\frac{1}{2}\left(-\frac{p_{t}^{2}f^{\prime}(r)}{f(r)^{2}}+\frac{p_{r}^{2}f^{\prime}(r)^{-1}}{f(r)^{-2}}-\tilde{C}p_{\phi}^{2}\right).\\ \end{split}\right.$ (10)
+Subsequently, the equations of motion for null particles are
+[…]
+Using Eq.(12), we can derive
+[…]
 where $\tilde{C}\equiv\frac{d}{dr}\frac{1}{C(r)}=-\frac{2\left(1-\frac{b_{2}}{r}\right)^{-\frac{2\alpha^{2}}{1+\alpha^{2}}}}{r^{3}}-\frac{2\alpha^{2}b_{2}\left(1-\frac{b_{2}}{r}\right)^{-1-\frac{2\alpha^{2}}{1+\alpha^{2}}}}{\left(1+\alpha^{2}\right)r^{4}}$.
 Setting $H=0$, we have
-$-\frac{p_{t}^{2}}{f(r)}+\frac{p_{r}^{2}}{f(r)^{-1}}+\frac{p_{\phi}^{2}}{C(r)}=0,$ (11)
+$\frac{p_{t}^{2}}{f(r)}+\frac{p_{r}^{2}}{f(r)^{-1}}+\frac{p_{\phi}^{2}}{C(r)}=0,$ (14)
 and it now follows that
 $\frac{dr}{d\phi}=\frac{\dot{r}}{\dot{\phi}}=\frac{C(r)}{f(r)^{-1}}\frac{p_{r}}{p_{\phi}}.$ (15)
-Setting $p_{t}=-\omega_{0}$, and using $p_{r}$, we get the relation how $r$ changes with $\phi$:
+Substituting the expressions for $p_{r}$ and $p_{t}=-\omega_{0}$, we finally obtain
 $\frac{dr}{d\phi}=\pm\frac{r^{2}}{C(r)^{1/2}f(r)^{-1/2}}\sqrt{\frac{\omega^{2}_{0}}{p^{2}_{\phi}}h(r)^{2}-1},$ (16)
 where $h(r)^{2}=\frac{C(r)}{f(r)}$ is defined.
-For a circular light orbit, the radial velocity and acceleration should be $\dot{r}=0$ and $\ddot{r}=0$ respectively, and hence, $p_{r}=0$. Eq.(11) then becomes
+For a circular photon orbit, the conditions $\dot{r}=0$ (vanishing radial velocity) and $\ddot{r}=0$ (zero radial acceleration) necessarily imply $p_{r}=0$, which reduces Eq.(14) to the following form
 $0=-\frac{w^{2}_{0}}{f(r)}+\frac{p^{2}_{\phi}}{C(r)}.$ (17)
-Since $\dot{p}_{r}=0$, Eq.(14) can be rewritten as
-$\dot{p}_{r}=0=\frac{\omega_{0}^{2}f^{\prime}(r)}{f(r)^{2}}+\tilde{C}p^{2}_{\phi}.$ (15)
-Using Eq.(14) and (15), we find
-$p^{2}_{\phi}=C(r)\frac{\omega^{2}_{0}}{f(r)},$ (16)
-and
-$p^{2}_{\phi}=-\frac{\omega_{0}^{2}f^{\prime}(r)}{\tilde{C}f(r)^{2}}.$ (17)
-The implication of subtracting Eqs. (16) and (17) give the information on how to find the radius of the photon sphere 176 ; 184
+Furthermore, Eq.(17) can be expressed as
+$\dot{p}_{r}=\frac{\omega_{0}^{2}f^{\prime}(r)}{f(r)^{2}}+\tilde{C}p^{2}_{\phi}=0.$ (18)
+Using Eq.(17) and (18), we obtain
+$p^{2}_{\phi}=C(r)\frac{\omega^{2}_{0}}{f(r)}=-\frac{\omega_{0}^{2}f^{\prime}(r)}{\tilde{C}f(r)^{2}},$ (19)
+Combining Eq.(18) and Eq.(19), we give analytic expression for the photon sphere radius of the BH 176 ; 184 (See more details in Appendix A).
 $0=\frac{d}{dr}h(r)^{2}.$ (20)
-Figure 2: The above four pictures plot the photon sphere radius under the different parameters $\alpha$, $\beta$ and $q$.
 Following Eq.(20), the analytical form of the photon sphere radius $r_{\text{ph}}$ is so complicated.
@@
 Then, for a given value of $q$, we can observe an increase in photon sphere radius as the $\alpha$ increases.
-Let us now determine the behavior of the shadow radius.
-For constructing the shadow, we assume that a static observer at radius coordinate $r_{\text{o}}$ sends light rays into the past.
-As can be seen from Fig.3, the angle $a$ between such a light ray and the radial direction is given by
-$\cot a=\left.\frac{\sqrt{g_{rr}}}{\sqrt{g_{\phi\phi}}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}=\left.\frac{\sqrt{B(r)}}{\sqrt{C(r)}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}$ (19)
-With the help of Eq.(12), we obtain
-$\cot^{2}a=\frac{h\left(r_{\mathrm{O}}\right)^{2}}{h(R)^{2}}-1.$ (20)
-By elementary trigonometry, we get 176
-$\sin^{2}a=\frac{h(R)^{2}}{h\left(r_{\mathrm{O}}\right)^{2}},$ (21)
-where $R$ means the minimum radius that the light ray approaches the center and then goes out.
-The boundary curve of the shadow corresponds to past-oriented light rays that asymptotically approach one of the unstable circular light orbits at radius $r_{\mathrm{ph}}$.
-Therefore we have to consider the limit $R\rightarrow r_{\mathrm{ph}}$ in Eq.(21) for getting the angular radius $a_{\text{sh }}$ of the shadow 184 ,
+Figure 2: The above four pictures plot the photon sphere radius under the different parameters $\alpha$, $\beta$ and $q$.
+Subsequently, the definition of the shadow radius must be established.
+The shadow radius behavior can be investigated by considering light rays traced backward in time from a static observer located at radial coordinate $r_{O}$.
+As shown in Fig.3, the angle $a$ formed between such light rays and the radial direction is expressed as
+$\cot a=\left.\frac{\sqrt{g_{rr}}}{\sqrt{g_{\phi\phi}}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}=\left.\frac{\sqrt{B(r)}}{\sqrt{C(r)}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}.$ (21)
+With the help of Eq.(15), we obtain 176
+[…]
+where $R$ means the minimum radius that the light ray approaches the center and then goes out (turning point of the trajectory).
+The shadow boundary is determined by past-oriented light rays asymptotically approaching the unstable circular photon orbits at radius $r_{\mathrm{ph}}$.
+Consequently, the angular radius $a_{\text{sh}}$ of the shadow is obtained by taking the limit $R\rightarrow r_{\mathrm{ph}}$ in Eq.(22) 184 .
 $\sin^{2}a_{\mathrm{sh}}=\frac{h\left(r_{\mathrm{ph}}\right)^{2}}{h\left(r_{\mathrm{O}}\right)^{2}}.$ (23)
@@
 $b_{\mathrm{c}}=h\left(r_{\mathrm{ph}}\right).$ (24)
-Therefore we can also write Eq.(22) as
+Therefore, the Eq.(23) can also be expressed as
 $\sin^{2}a_{\mathrm{sh}}=\frac{b_{\mathrm{c}}^{2}}{h\left(r_{\mathrm{O}}\right)^{2}},\quad\text{ or }\quad\sin^{2}a_{\mathrm{sh}}=\frac{b_{\mathrm{c}}^{2}A\left(r_{\mathrm{O}}\right)}{C\left(r_{\mathrm{O}}\right)}.$ (25)
-For a general static spherically symmetric asymptotically flat spacetime, Ref.176 demonstrates that when the observer’s radial coordinate $r_{o}$ is sufficiently large, the black hole shadow radius approaches $b_{c}$.
+For a general static spherically symmetric asymptotically flat spacetime, Ref.176 demonstrates that when the observer’s radial coordinate $r_{O}$ is sufficiently large, the BH shadow radius approaches $b_{c}$.
 This radius is determined by the location of the photon sphere and the specific form of the metric function $f(r)$.
 In our paper, we will obey this consideration.
-Figure 3: Example of calculation of light ray emitted from the observer’s position (small disk) into the past under an angle $a$.
+Figure 3: Example of calculation of light ray emitted from the observer’s position into the past under an angle $a$.
 The BH horizon and the photon sphere are shown; $r_{\mathrm{ph}}$ is the photon sphere radius.
-The trajectory is calculated in the Asymptotically flat spacetime, $R$ denotes the radius coordinate at the point of closest approach.
 The picture is indicated in Ref.176 .
 III.1 Constraints on parameters with the EHT observations of M87* and Sgr A*
-In this subsection, we find constraints to the coupling parameters $\alpha$ and $\beta$ using the observation data provided by the EHT for M87* and Sgr A*.
-We only focused on the non-rotating case since the rotation parameter of Sgr A* is small enough to have considerable deviation to the shadow radius 110 .
-Furthermore, it has also been concluded for M87* that it is difficult to distinguish between Kerr and dilaton BHs based on BH shadow images alone using general-relativistic magnetohydrodynamical simulations and radiative-transfer calculations to generate synthetic shadow images in comparison to the present observation from the EHT 111 .
-Also remarked in Ref.112 that the shadow size of M87* lies within the range of $3\sqrt{3}(1\pm 0.17)M$, whether their model is spherically symmetric or axisymmetric.
-[…]
-$d_{\mathrm{sh}}=\frac{D\theta}{M}$ (25)
-Figure 4: These plots are showing the constraints for different coupling parameters $\alpha$ and $\beta$.
-Hence, the theoretical shadow diameter, however, can be obtained via
+We will obtain constraints on the coupling parameters $\alpha$ and $\beta$ using EHT data for M87 and Sgr A, focusing on the non-rotating case. This simplification is justified because
+  • M87: General-relativistic magnetohydrodynamic simulations indicate that current EHT observations cannot reliably distinguish between Kerr and dilaton BHs based on shadow morphology alone 111 .
+The observed shadow size is consistent with $3\sqrt{3}(1\pm 0.17)M$, regardless of spherical or axisymmetric models 112 .
+  • Sgr A: The rotation parameter is sufficiently small to negligibly affect the shadow radius 110 .
+Based on observational data, the shadow of M87* exhibits an angular diameter of $42\pm 3\ \mathrm{\rm\mu as}$, corresponding to a distance of $16.8$ Mpc from Earth and a central BH mass of $(6.5\pm 0.90)\times 10^{9}\ M_{\odot}$ 113 .
+In parallel, the EHT’s recent findings for Sgr A* report a shadow size of $48.7\pm 7\ \mathrm{\mu as}$, with the source located at a distance of $8277\pm 33$ pc and hosting a supermassive BH of mass $(4.3\pm 0.013)\times 10^{6}\ M_{\odot}$ 114 ; 115 .
+The shadow diameter in mass units can be calculated from the derived BH parameters using the following expression 116
+$d_{\mathrm{sh}}=\frac{D\theta}{M}.$ (26)
+Hence, the theoretical shadow diameter can be obtained via
 $d_{\text{sh }}^{\text{theo }}=2R_{\text{sh}}.$ (27)
 Therefore, by using the above expression, we get the diameter of the shadow of M87* and Sgr A*
-$\left\{\begin{split}&d_{\mathrm{sh}}^{\mathrm{M}87^{*}}=(11\pm 1.5)M,\\ &d_{\mathrm{sh}}^{\mathrm{Sgr}.\mathrm{A}^{*}}=(9.5\pm 1.4)M.\\ \end{split}\right.$ (27)
-M87*$/\alpha$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$ M87*/$\beta$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$
+[…]
+Our constraints on the model parameters are shown in Fig.4. The corresponding numerical constraints are further presented in Tab.I (for M87*) and Tab.II(for Sgr A*).
+Figure 4: These plots are showing the constraints for different coupling parameters $\alpha$ and $\beta$.
+[…]
 charge $Q$ upper lower upper lower charge $Q$ upper lower upper lower
@@
 TABLE I. The table lists the values of $\alpha$ and $\beta$ at the $1\sigma$ and $2\sigma$ confidence levels for M87*.
-Sgr A*$/\alpha$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$ Sgr A*/$\beta$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$
+[…]
 charge $Q$ upper lower upper lower charge $Q$ upper lower upper lower
@@
 The variation of the diameter of the shadow image with coupling parameter $\alpha$ and $\beta$ for M87* and for Sgr A* is shown in Fig.4, showing uncertainties at $1\sigma$ and $2\sigma$ levels.
-The numerical values for the upper or lower bounds in $\beta$ and $\alpha$ is found in Table I and Table II.
+The numerical values for the upper or lower bounds in $\beta$ and $\alpha$ is found in Tab.I and Tab.II.
 However, it turns out that the data for M87* gives a better constraint for the coupling parameter $\beta$.
@@
 III.2 Light bending: direct emission, lensing ring and photon ring
-To gain a more comprehensive understanding of the appearance of EMS BH, we analyze its photon rings, and lensing rings in the presence of an optically bright accretion disk 177 ; 178 ; 179 .
-We begin by investigating the trajectory of light rays propagating in the vicinity of the EMS BH.
-To facilitate this analysis, we introduce the transformation $u=1/r$, which reformulates the orbit equation as follows
-$(\frac{du}{d\phi})^{2}=G(u),$ (28)
-where
+To systematically investigate the observational characteristics of EMS BH, we examine their photon ring and lensing ring structures under the assumption of an optically thick accretion disk 177 ; 178 ; 179 .
+Our analysis begins with studying null geodesics in the EMS BH.
+For computational convenience, we employ the coordinate transformation $u\equiv 1/r$, which yields the compact orbital equation
+$\left(\frac{du}{d\phi}\right)^{2}=G(u),$ (29)
+where the function $G(u)$ takes the explicit form
 $\sqrt{G(u)}=\frac{1}{C(1/u)^{1/2}f(1/u)^{-1/2}u^{2}}\sqrt{\frac{\omega^{2}_{0}}{p^{2}_{\phi}}h(1/u)^{2}-1}.$ (30)
-For a static and spherically symmetric spacetime, we consider the motion of a particle confined to the equatorial plane. From the Lagrangian, two conserved quantities can be obtained from Eq. (10).
-The impact parameter is defined as $b\equiv\frac{p_{t}}{p_{\phi}}$.
-When $b>b_{c}$, a light ray originating from infinity approaches the black hole, reaches a closest approach, and then escapes back to infinity.
-For $b<b_{c}$, the light ray inevitably falls into the black hole.
-In the critical case $b=b_{c}$, the light ray asymptotically orbits the black hole at the photon sphere radius $r_{\mathrm{ph}}$.
-For $b>b_{c}$, the turning point of the trajectory corresponds to the smallest positive real root of $G(u)=0$, denoted as $u_{m}$.
-According to Eq.
-(28), the total change in the azimuthal angle $\phi$ along a trajectory with a given impact parameter $b$ can be expressed as
-$\phi=2\int_{0}^{u_{m}}\frac{du}{\sqrt{G(u)}},\quad b>b_{c}.$ (30)
-For $b<b_{c}$, we only focus on the trajectory outside the horizon, so the total change of azimuthal angle $\phi$ is obtained by
-$\phi=\int_{0}^{u_{0}}\frac{du}{\sqrt{G(u)}},\quad b<b_{c}$ (31)
-where $u_{0}=1/r_{+}$.
-Figure 5: The behavior of photon trajectories around the Schwarzschild, RN, and EMS black holes as a function of the impact parameter b.
+Similarly, we restrict our analysis to equatorial motion.
+The system’s Lagrangian formalism provides two conserved quantities, as derived in Eq.(13).
+The critical impact parameter $b_{c}$ separates distinct null geodesic behaviors
+  • Case 1 ($b>b_{c}$): Photons from asymptotic infinity are gravitationally scattered, reaching a minimum coordinate distance $r_{\text{min}}$ before returning to infinity.
+  • Case 2 ($b<b_{c}$): Photons become captured by the BH, demonstrating the existence of a photon capture cross-section.
+  • Case 3 ($b=b_{c}$): Photons spiral asymptotically toward the photon sphere at $r_{\text{ph}}$, defining the boundary of BH shadow.
+For $b>b_{c}$, the turning point of the trajectory corresponds to the smallest positive real root of $G(u)=0$, denoted as $u_{\rm m}=1/r_{\rm min}$ with $r_{\rm min}$ being the light ray’s radial minimal distance from its trajectory to the BH.
+Hence, the total change in the azimuthal angle $\phi$ outside the horizon of the trajectory with a given impact parameter $b$ can be expressed as 178 ; 179
+$\phi=2\int_{0}^{u_{\rm m}}\frac{du}{\sqrt{G(u)}},\quad b>b_{c}.$ (31)
+For $b<b_{c}$, the total change of angle $\phi$ is given by
+$\phi=\int_{0}^{u_{0}}\frac{du}{\sqrt{G(u)}},\quad b<b_{c},$ (32)
+where $u_{0}\equiv 1/r_{+}$.
+Figure 5: The behavior of photon trajectories around the Schwarzschild, RN, and EMS BHs as a function of the impact parameter $b$.
 In the upper panel, we present the total number of orbits, defined as $n=\phi/2\pi$.
-The trajectories are categorized based on $n$, where direct emission $n<3/4$ is shown in black, lensed trajectories 3/$43/4<n<5/$ 4 in yellow, and photon ring trajectories $n>5/4$ in red.
+The trajectories are categorized based on $n$, where direct emission $n<3/4$ is shown in black, lensed trajectories $3/4<n<5/4$ in yellow, and photon ring trajectories $n>5/4$ in red.
 The lower panel displays selected photon trajectories in Euclidean polar coordinates $(r,\phi)$.
@@
 For the three cases studied, we set $\alpha=0,\beta=0,q=0$ for the Schwarzschild BH (first column), $\alpha=0,\beta=0,q=0.5$ for the RN BH (second column), and $\alpha=0.5,\beta=0.8,q=0.5$ for the EMS BH (third column).
-To analyze the observational signatures of radiation emitted from the vicinity of a BH accretion disk.
-The Ref.117 divide trajectories into direct, lensed and photon rings.
-One define the total number of orbits $n=\frac{\phi}{2\pi}$ which is obviously a function of impact parameter $b$.
-We denote the solution of
-$n(b)=\frac{2m-1}{4},\quad m=1,2,3,\cdots$ (32)
-by $b_{m}^{\pm}$. Note that $b_{m}^{-}<b_{c}$ and $b_{m}^{+}>b_{c}$. Then we can classify all trajectories as follows
+Following the trajectory classification established in Refs.
+120 ; 178 where photon orbits are categorized into direct, lensed, and photon ring components, we quantify the orbital winding through the turn number $n(b)\equiv\phi/2\pi$.
+This quantity exhibits discrete quantization as demonstrated in Refs.
+178 ; 179 , with the functional dependence on impact parameter $b$ given by
+$n(b)=\frac{2m-1}{4},\quad m\in\mathbb{Z}^{+}.$ (33)
+The critical impact parameters $b_{m}^{\pm}$ demarcate these orbital families, where $b_{m}^{-}$ (satisfying $b_{m}^{-}<b_{c}$) and $b_{m}^{+}$ (satisfying $b_{m}^{+}>b_{c}$) correspond respectively to the minimal and maximal solutions for each $m$-th order orbit, with $b_{c}$ denoting the photon sphere critical parameter.
+Then the rays can be classified as follows
   • Direct: $\frac{1}{4}<n<\frac{3}{4}\Rightarrow b\in\left(b_{1}^{-},b_{2}^{-}\right)\cup\left(b_{2}^{+},\infty\right)$
@@
   • Photon ring: $n>\frac{5}{4}\Rightarrow b\in\left(b_{3}^{-},b_{3}^{+}\right)$
-The physical picture of this classification is clear from the trajectory plots in Fig.5.
-Assuming light rays emit from north pole direction (far right of the trajectory plots), trajectories whose number of orbits 1/$41/4<n<3/$ 4 will intersect the equatorial plane only once.
-Trajectories whose number of orbits 3/$43/4<n<5/$ 4 will intersect the equatorial plane twice.
-Trajectories whose number of orbits $n>5/4$ will intersect the equatorial plane at least 3 times.
+The behaviors of photons in different BHs are plotted in Fig.5.
+The up panel of Fig.5 depicts the total number of orbits.
+And the down panel depicts the trajectories of light rays surrounding the BHs, such as ”direct” will intersect the equatorial plane only once.
 Parameter $\alpha=0,\beta=0,q=0$ $\alpha=0,\beta=0,q=0.5$ $\alpha=0.5,\beta=0.8,q=0.5$
@@
 $n<3/4$ $b>6.16757$ $b>5.97448$ $b>5.81006$
-Lensing Ring $5.187815.01514<b<$ 5.18781 $4.957934.77294<b<$ 4.95793 $4.753924.55452<b<$ 4.75392
-3/$43/4<n<5/$ 4 $6.167575.22794<b<$ 6.16757 $5.974485.0039<b<$ 5.97448 $5.810064.80685<b<$ 5.81006
+Lensing Ring $5.01514<b<5.18781$ $4.77294<b<4.95793$ $4.55452<b<4.75392$
+$3/4<n<5/4$ $5.22794<b<6.16757$ $5.0039<b<5.97448$ $4.80685<b<5.81006$
 Photon Ring
-$n>5/4$ $5.227945.18781<b<$ 5.22794 $5.00394.95793<b<$ 5.0039 $4.806854.75392<b<$ 4.80685
+$n>5/4$ $5.18781<b<5.22794$ $4.95793<b<5.0039$ $4.75392<b<4.80685$
 TABLE III. The region of direct emission, lensing ring, and photon ring for the different parameter $\alpha,\beta,q$.
@@
 III.3 Transfer functions and observed specific intensities
-We analyze the emitted intensity from the EMS BH, assuming that the disk radiates isotropically in the rest frame of a static observer.
-According to Liouville’s theorem, the quantity $I_{\nu}^{\mathrm{em}}/\nu_{e}^{3}$ remains conserved along the trajectory of the light ray.
-Consequently, the observed intensity can be expressed as
+In this subsection, we will derive the specific intensity $I_{\nu}^{\mathrm{em}}$ radiated by the accretion disk, assuming local isotropy of radiation in the rest frame of static observers.
+The observed intensity can be expressed as 177 ; 178 ; 179
 $I_{\nu^{\prime}}^{\mathrm{obs}}=g^{3}I_{\nu}^{\mathrm{em}},$ (34)
-[…]
+where $g=\sqrt{f(r)}$ and $I_{\nu^{\prime}}^{obs}$ are the observed intensity at the frequency $\nu^{\prime}$.
+By integrating over all frequencies, we obtain the total emitted intensity as $I^{\mathrm{em}}=\int I_{\nu}^{\mathrm{em}}d\nu$.
+Thus, the corresponding observed frequency can be written as
 $I^{\mathrm{obs}}=g^{4}I^{\mathrm{em}}.$ (35)
-$I_{\mathrm{em}}$ here is the total emitted specific intensity from the accretion disk. Therefore, the total intensity received by the observer will be
+$I^{\mathrm{em}}$ denotes the integrated specific intensity radiating from the entire accretion disk. Hence, the total received intensity is
 $I^{\mathrm{obs}}(b)=\left.\sum_{n}g^{4}I^{\mathrm{em}}(r)\right|_{r=r_{m}(b)},$ (36)
-where $r_{m}(b)$ is the $m$ intersection outside the horizon in the equatorial plane, which we call the transfer function.
-The transfer function establishes the connection between the radial coordinates and the photon’s impact parameter.
-It is important to note that our analysis does not account for light absorption or reflection by the accretion disk, nor does it consider the attenuation of light intensity due to environmental effects.
-Instead, we work within an idealized model.
-We denote the solution of the orbit equation by $u(\phi,b)$ and focus on the first three transfer functions, which can be obtained by
-$\left\{\begin{split}&r_{1}(b)=\frac{1}{u\left(\frac{\pi}{2},b\right)},&b\in\left(b_{1}^{-},\infty\right)\\ &r_{2}(b)=\frac{1}{u\left(\frac{3\pi}{2},b\right)},&b\in\left(b_{2}^{-},b_{2}^{+}\right)\\ &r_{3}(b)=\frac{1}{u\left(\frac{5\pi}{2},b\right)},&b\in\left(b_{3}^{-},b_{3}^{+}\right)\end{split}\right.$ (36)
-As described in Ref.120 , the first transfer function corresponds to the “direct image” of the disk, which primarily reflects the redshift of the source profile.
-The second transfer function produces a highly demagnified image of the disk’s far side, known as the “lensing ring”.
-The third transfer function yields an extremely demagnified image of the near side of the disk, referred to as the “photon ring”.
-Images generated by subsequent transfer functions are so severely demagnified that they can be effectively disregarded.
+where $r_{m}(b)$ is transfer function (It represents the $m$ intersection points with the equatorial plane).
+The transfer function fundamentally connects radial coordinates with photon impact parameters, we emphasize that our analysis operates under two key idealizations: (1) no light absorption or reflection by the disk, and (2) no environmental intensity attenuation.
+Therefore, the first three transfer functions, which can be obtained by 177 ; 178 ; 179
+$\left\{\begin{split}&r_{1}(b)=\frac{1}{u\left(\frac{\pi}{2},b\right)},&b\in\left(b_{1}^{-},\infty\right),\\ &r_{2}(b)=\frac{1}{u\left(\frac{3\pi}{2},b\right)},&b\in\left(b_{2}^{-},b_{2}^{+}\right),\\ &r_{3}(b)=\frac{1}{u\left(\frac{5\pi}{2},b\right)},&b\in\left(b_{3}^{-},b_{3}^{+}\right).\end{split}\right$.
+(37)
+As established in Ref.120 , three distinct observational features emerge:(1) The first transfer function corresponds to the direct disk image;(2) The second forms a significantly demagnified lensing ring from the disk’s far side;(3) The third creates the extremely faint photon ring from the near side.
+The corresponding graphical representation is presented in the figure below.
 Figure 6: The first three transfer functions in BHs for different values of $\alpha,\beta,q$.
@@
 Here, the y-axis is the transfer function $r_{m}(b)$ and the x-axis is the impact parameter b.
-In Fig.6, the black dots $(m=1)$ represent the transfer function for direct emission and exhibit an almost constant slope, indicating a redshifted source profile.
-The yellow dots $(m=2)$ correspond to the lensing ring, where the slope remains small near $b_{c}$ but increases rapidly with $b$, leading to significant demagnification of the image from the far side of the accretion disk.
-The red dots $(m=3)$ represent the photon ring, where the slope approaches infinity, implying extreme demagnification of the disk’s near-side image.
-Consequently, the dominant contribution to the observed flux arises from direct emission, while higher-order images $(m>3)$ contribute negligibly and can be disregarded.
-All three black holes exhibit these characteristics.
+As shown in Fig.6, the black points $(m=1)$ correspond to the direct emission transfer function and exhibit an approximately constant slope, reflecting the redshifted emission profile from the disk.
+The yellow points $(m=2)$, associated with the lensing ring, display a rapidly increasing slope with growing impact parameter $b$, indicating that the far-side image is significantly demagnified.
+The red points $(m=3)$ denote the photon ring, where the slope becomes nearly infinite, resulting in extreme demagnification of the near-side image.
+Thus, the observed flux is primarily governed by the direct emission, with higher-order contributions $(m>3)$ being negligible.
+All three BHs exhibit these characteristics.
 The critical impact parameters are $b_{c}=5.19615$ for the Schwarzschild BH, $b_{c}=4.96791$ for the RN BH, and $b_{c}=4.76598$ for the EMS BH.
 This result indicates that the Schwarzschild BH has the largest critical impact parameter, corresponding to the largest shadow radius, while an increase in the model parameters leads to a reduction in the BH’s shadow size.
+Additionally, we not only learn the aforementioned characteristics but can also quantify both the morphology (width) and intensity changes of the three ring types.
+First, we discuss how scalar couplings affect the photon ring.
+As shown in Table III, we have following calculations.
+In Schwarzschild BH ($q=0,\alpha=0,\beta=0$), we have $\Delta b_{\text{ph}}=b_{3}^{+}-b_{3}^{-}=5.22794-5.18781=0.04013$.
+In EMS BH ($\alpha=0.5,\beta=0.8,q=0.5$), we have $\Delta b_{\text{ph}}=b_{3}^{+}-b_{3}^{-}=4.80685-4.75392=0.05293$.
+The EMS BH shows a net increase of $32\%$ in width.
+A possible explanation is: the scalar field flattens the effective potential near $r_{\text{ph}}$ (due to $C(r)=r^{2}(1-b_{2}/r)^{\frac{2\alpha^{2}}{1+\alpha^{2}}}$), allowing more near-critical orbits.
+For lensing ring adjustment, we now properly treat the two disjoint intervals ($b\in(b_{2}^{-},b_{3}^{-})$ and $b\in(b_{3}^{+},b_{2}^{+})$).
+In Schwarzschild BH ($q=0,\alpha=0,\beta=0$), we have $\Delta b_{\text{lens1}}=b_{3}^{-}-b_{2}^{-}=5.18781-5.01514=0.17267$, $\Delta b_{\text{lens2}}=b_{2}^{+}-b_{3}^{+}=6.16757-5.22794=0.93963$.
+Total $\Delta b_{\text{lens}}=1.1123$.
+In EMS BH ($\alpha=0.5,\beta=0.8,q=0.5$), we have $\Delta b_{\text{lens1}}=b_{3}^{-}-b_{2}^{-}=4.75392-4.55452=0.1994$, $\Delta b_{\text{lens2}}=b_{2}^{+}-b_{3}^{+}=5.81006-4.80685=1.00321$.
+Total $\Delta b_{\text{lens}}=1.20261$.
+Compared to a Schwarzschild BH, the EMS BH shows a net increase of $8.1\%$.
+While the total lensing ring width increases modestly, the inner sub-interval ($\Delta b_{\text{lens1}}$) shows a $15.5\%$ increase due to stronger light bending near the photon sphere.
+Finally we consider the direct emission, increasing $\alpha$ leads to a larger $b_{c}$, whereas increasing $\beta$ decreases $b_{c}$.
+From Table III, the boundaries of the direct emission region exhibit a systematic shift along the impact-parameter axis.
 III.4 Observational features of direct emission, photon and lensing rings
 As discussed in the previous section, for an observer at infinity, the specific intensity depends solely on the radial coordinate $r$.
-In our analysis, we consider three toy models for the emissivity profile $I^{\mathrm{em}}$.
-With the transfer function determined, we now consider a specific emission profile.
-As the first model, we examine a scenario in which the emission is sharply concentrated at the innermost stable circular orbit (ISCO) given by
+In our analysis, we consider the following three toy models for the emissivity profile $I^{\mathrm{em}}$177 ; 178 ; 179 ,
 $I_{1}^{\mathrm{em}}(r)=\begin{cases}\left(\frac{1}{r-\left(r_{\mathrm{isco}}-1\right)}\right)^{2},&r>r_{\mathrm{isco}}\\ 0,&r\leq r_{\mathrm{isco}}\end{cases}$ (38)
-Second model, we consider the emission is sharply peaked at the photon sphere, and it ends abruptly at $r=r_{\text{ph}}$ while quickly decaying to zero, such as
 $I_{2}^{\mathrm{em}}(r)=\begin{cases}\left(\frac{1}{r-\left(r_{\mathrm{ph}}-1\right)}\right)^{3},&r>r_{\mathrm{ph}}\\ 0,&r\leq r_{\mathrm{ph}}\end{cases}$ (39)
-Final model, we consider an emission decaying gradually from the horizon to the ISCO, such as
 $I_{3}^{\mathrm{em}}(r)=\begin{cases}\frac{1-\arctan\left(r-\left(r_{\text{isco }}-1\right)\right)}{1-\arctan\left(r_{\mathrm{ph}}\right)},&r>r_{\mathrm{+}}\\ 0,&r\leq r_{\mathrm{+}}\end{cases}$ (40)
-Each of these three models exhibits distinct characteristics.
-In particular, the second model features a rapid decay rate, whereas the third model exhibits a much slower decay.
-In the third model, the emission originates directly from the event horizon, while in the second model, it begins at the photon sphere, and in the first model, it is confined to the ISCO.
-Although these models represent highly idealized scenarios, they provide valuable qualitative insights into the behavior of photons around the BH.
+where $r_{\text{isco}},r_{\text{ph }}$ and $r_{\mathrm{+}}$ is the innermost stable orbit, photon sphere and event horizon respectively.
+These three models each exhibit distinct characteristics.
+The following figures present the results for three distinct types of BHs.
 Figure 7: The observational appearance of the thin disk with different emission profiles for $\alpha=0,\beta=0,q=0$ is shown from a face-on perspective.
@@
 This figure presents the results of the three models corresponding to EMS BH.
-In the Fig.7, 8, and 9, we have shown the observed appearance of the thin accretion disk with the observed intensity and the impact factor corresponding to the three models for different coupling constant parameters respectively.
-In all the figures, we see that for the first model (top row), the emission intensity (top-first) has a peak near the critical impact parameter $r\sim b_{c}$ and then it decreases as radial distance increases and becomes zero.
-In this case, the photon sphere lies in the interior region of the emission part of the disk.
-As can be readily seen from the middle panel (top-second), the direct image of the disk looks very similar to the emission profile, although its abrupt end occurs at different place due to gravitational lensing.
-However, we noticed that the peak of the photon and lensing rings are not only smaller than the direct emission but also have a narrow observational area.
-Hence, we observed that the observed intensity has a huge contribution coming from the direct emission, a small contribution from the lensing rings, and very little contribution from the photon rings.
-As is evident from the right panel (2D shadow image, top-third) where the lensing ring emission appears as a tiny ring inside the direct image and one can barely see the photon ring emission in the right panel if one zooms in.
-For the second model (middle row), the intensity of the emission has a peak at photon sphere $r\sim r_{\mathrm{ph}}$ and then decreases as radial distance increases (middle-first).
-In this case, the observed intensity profile (middle-second) peaks due to the direct emission, and then it shows an attenuation with increasing $r$.
-We observed that the photon and lensing rings, which improves the total intensity of this particular area, and therefore, we get a new peak due to the photon ring, lensing ring, and direct emission.
-However, the photon and lensing rings are highly demagnetized and have a very narrow area in the observed intensity.
-Therefore, we still have a dominant contribution coming from the direct emission in the observed intensity which can be seen in the $2\mathrm{D}$ shadow image (middle-third).
-For the third model (bottom row), the peak in the intensity starts from the horizon ($r_{+}$) and then decreases with increasing $r$ (bottom-first).
-In this case, redshift effects noticeably decrease the observed flux.
-However, the most important difference from the top two row for our considerations is that the lensing ring and photon ring emission are now superimposed on the direct emission.
-However, as discussed in the last section, the photon and lensing rings are highly demagnetized and have a very narrow area in the observed intensity.
-Therefore, we still have a dominant contribution coming from the direct emission in the observed intensity which can be seen in the 2D shadow image (bottom-third).
+Figures 7, 8, and 9 display the observed emission from thin accretion disks, intensity distributions, and impact parameter mappings for three BH models under different coupling parameters.
+In each case, the top row corresponds to the first model.
+The intensity profile (top-left panel) peaks near the critical impact parameter $b_{c}$ and gradually decreases outward.
+The direct image (top-middle panel) closely follows the intrinsic emission but is affected by gravitational lensing, leading to a sharp boundary and image distortion.
+Notably, the photon and lensing rings contribute marginally to the overall intensity due to their limited angular extent and lower peak brightness.
+The shadow image (top-right panel) reveals the lensing ring as a narrow, bright ring just inside the direct image, while the photon ring appears extremely faint and confined to a narrow region, visible only upon magnification.
+These results highlight the dominant role of direct emission in observed disk images.
+In the second model (middle row), the emission intensity peaks near the photon sphere ($r\sim r_{\mathrm{ph}}$) and gradually decreases outward (middle-left panel).
+The observed intensity profile (middle-center) similarly shows a peak dominated by direct emission from this region, followed by a steady decline.
+Although the photon and lensing rings contribute modestly, their overlap with direct emission slightly enhances the total signal.
+Nonetheless, due to strong gravitational redshift and depolarization, their brightness remains faint and confined to a narrow angle.
+As shown in the shadow image (middle-right panel), direct emission remains the dominant observable feature.
+In the third model (bottom row), the emission intensity peaks near the event horizon ($r_{+}$) and gradually declines with increasing radius (bottom-left panel).
+Due to strong gravitational redshift, much of this radiation is suppressed before reaching a distant observer.
+Unlike the previous models, the photon and lensing ring emissions overlay the direct radiation; however, their contributions remain minimal because of severe demagnification and narrow angular extent.
+Consequently, the observed intensity is still primarily determined by direct emission, as clearly seen in the shadow image (bottom-right panel).
 Although Fig.7, 8 and 9 show only a few highly idealized cases of thin disk emission near a BH (viewed face on), it illustrates two key points that we believe will hold quite generally for optically thin disk emission: (1) The emission is dominated by the direct emission, with the lensing ring emission providing only a small contribution to the total flux and the photon ring providing a negligible contribution in all cases.
 (2) For all three models, as we increase the coupling constant $\alpha,\beta$ and $q$, the observed intensity decreases, and it is much less than what we get for the Schwarzschild BH.
+III.5 Weak gravitational lensing and magnification of lensed image
+Let us now examine the brightness of the image in the weak field limit by using the light’s deflection angle around EMS BH.
+We will find the deflection angle by using the Gauss-Bonnet theorem (GBT), originally stated as 667 ; 668
+$\iint_{D}KdS+\sum_{a=1}^{N}\int_{\partial D_{a}}\kappa_{\mathrm{g}}d\ell+\sum_{a=1}^{N}\theta_{a}=2\pi\chi(D).$ (41)
+Consider a freely orientable $2D$ curved surface $S$ with an infinitesimal area element $dS$, which defines a domain $D$ with a Gaussian curvature $K$.
+The boundary of $D$ is given by $N$ components, $\partial D_{\mathrm{a}}$ (where $\mathrm{a}=1,2,\ldots,N$), along which the geodesic curvature $\kappa_{\mathrm{g}}$ is integrated over the path element $d\ell$ following a positive convention.
+The jump angles are denoted by $\theta_{\mathrm{a}}$.
+The Euler characteristic of the domain, $\chi(D)$, is equal to 1 in this case since $D$ is a non-singular region.
+It was shown by Ref.666 that in spherically symmetric spacetime admitting asymptotic flatness, the deflection angle can be written as
+$\hat{\alpha}=-\iint_{D}K\,dS.$ (42)
+Substituting the EMS solution into the above expression and performing the weak-field approximation yields
+$\hat{\alpha}\approx\frac{4M}{b}-\frac{3\pi Q^{2}\left(1-\alpha^{2}+\beta\right)}{4b^{2}}.$ (43)
+In the context of gravitational lensing, let us consider the following lens equation, expressed in terms of the light angles $\hat{\alpha}$, $\theta$, and $\eta$ Morozova13 ; Bozza2008lens ; Babar21a ; Al_Badawi_2024
+$\theta D_{\mathrm{s}}=\eta D_{\mathrm{s}}+\hat{\alpha}D_{\mathrm{ds}}\,$, (44)
+where $D_{\mathrm{s}}$ is the distance between the source and observer, $D_{\mathrm{ds}}$ is the distance between the source and lens, and the angular positions of the image and the source are denoted by $\theta$ and $\eta$, respectively.
+From Eq.
+(44), the source’s angular position can be expressed as
+$\eta=\theta-\frac{D_{\mathrm{ds}}}{D_{\mathrm{s}}}\frac{\hat{\alpha}(\theta)}{D_{\mathrm{d}}}\,$, (45)
+where $D_{\mathrm{d}}$ is the distance between the lens and observer.
+We note the use of the reduced deflection angle $\hat{\alpha}(\theta)=\xi(\theta)/b$, where $b\equiv D_{d}\theta$ is the impact parameter Bozza2008lens .
+The angular part, known as the Einstein radius $\theta_{E}$, is derived from the spacetime geometry between the source images and corresponds to the radius of the Einstein ring, provided the image forms a ring-like shape.
+It is given by Morozova13
+$\theta_{E}=\sqrt{2R_{s}\frac{D_{ds}}{D_{d}D_{s}}}\,$, (46)
+with $R_{s}=D_{d}\theta_{E}$ as the radius of the Einstein ring.
+Next, we investigate the magnification of brightness, defined as the ratio of the total brightness of the lensed image ($I_{\mathrm{tot}}$) to the unlensed source brightness ($I_{*}$).
+The total magnification is given by Babar21a ; Atamurotov2022
+$\mu_{\mathrm{tot}}=\frac{I_{\mathrm{tot}}}{I_{*}}=\sum_{k}\bigg|\frac{\theta_{k}}{\beta}\frac{d\theta_{k}}{d\beta}\bigg|,\quad k=1,2,\dots,j\,.$ (47)
+For the specific case of two images, the total magnification simplifies to
+$\mu_{\mathrm{tot}}=\frac{x^{2}+2}{x\sqrt{x^{2}+4}}\,$, (48)
+where $x=\eta/\theta_{E}$ is a dimensionless parameter.
+We conducted a numerical analysis of the total magnification for different values of the parameters $\alpha$, $\beta$, and $q$ for EMS BH (Fig  10).
+Our findings show that for a fixed value of $q$, the total magnification increases with increasing $\alpha$ but decreases with increasing $\beta$.
+Furthermore, the total magnification decreases as the BH charge increases.
+We also observed a downward shift in the magnification curves for increasing values of $\beta$ when $\alpha=0$.
+Conversely, for $\beta=0$, we obtain the opposite trend compared to the $\alpha=0$ case.
+Figure 10: Total magnification $\mu_{tot}$ plotted as a function of the different parameters $\alpha$, $\beta$ and $q$.
 IV Conclusions
@@
 In other words, an increase in the model parameters leads to a reduction in the disk’s radiation intensity.
+Fianlly, when we consider weak gravitational lensing in EMS BH, our results indicate that for fixed $q$, the total magnification grows with $\alpha$ but decreases with $\beta$ and with increasing BH charge.
+Moreover, $\beta$ shifts the curves downward when $\alpha=0$, while for $\beta=0$ the trend reverses.
 Acknowledgements.
@@
 ZDSYS20190902092905285), and Center for Computational Science and Engineering at Southern University of Science and Technology.
+Appendix A Derivations for $r_{ph}$
+For a general static, spherically symmetric spacetime, the photon sphere radius $r_{\mathrm{ph}}$ can be determined by the extremum condition of the effective potential for null geodesics.
+Using the metric
+$ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+C(r)\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right),$ (49)
+the condition for the photon sphere reads
+$\frac{d}{dr}\left(\frac{C(r)}{f(r)}\right)=0,$ (50)
+or equivalently:
+$f^{\prime}(r)\,C(r)-f(r)\,C^{\prime}(r)=0.$ (51)
+In our EMS black hole model, the metric functions are
+$\begin{cases}f(r)=\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{\frac{1-\alpha^{2}}{1+\alpha^{2}}}+\frac{\beta Q^{2}}{C(r)},\\ C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}},\end{cases}$ (52)
+with
+$\begin{cases}b_{1}=\left(1+\sqrt{1-q^{2}(1-\alpha^{2})}\right)M,\quad\\ b_{2}=\frac{1+\alpha^{2}}{1-\alpha^{2}}\left[1-\sqrt{1-q^{2}(1-\alpha^{2})}\right]M.\end{cases}$ (53)
+Now, for simplicity, we define
+$p=\frac{1-\alpha^{2}}{1+\alpha^{2}},\quad s=\frac{2\alpha^{2}}{1+\alpha^{2}},$ (54)
+then we have
+$\begin{cases}C^{\prime}(r)&=2r\left(1-\frac{b_{2}}{r}\right)^{s}+sb_{2}\left(1-\frac{b_{2}}{r}\right)^{s-1},\\ f^{\prime}(r)&=\frac{b_{1}}{r^{2}}\left(1-\frac{b_{2}}{r}\right)^{p}+\left(1-\frac{b_{1}}{r}\right)p\left(1-\frac{b_{2}}{r}\right)^{p-1}\frac{b_{2}}{r^{2}}-\beta Q^{2}\frac{C^{\prime}(r)}{C(r)^{2}}.\end{cases}$ (55)
+Substituting $f(r),f^{\prime}(r),C(r)$ and $C^{\prime}(r)$ into the photon sphere condition
+$f^{\prime}(r)\,C(r)-f(r)\,C^{\prime}(r)=0,$ (56)
+we obtain the explicit nonlinear equation for $r_{\mathrm{ph}}$ in terms of the model parameters $\alpha,\beta,Q,M$,
+$0=$ ${\left[\frac{b_{1}}{r^{2}}\left(1-\frac{b_{2}}{r}\right)^{p}+\left(1-\frac{b_{1}}{r}\right)p\left(1-\frac{b_{2}}{r}\right)^{p-1}\frac{b_{2}}{r^{2}}-\beta Q^{2}\frac{2r\left(1-\frac{b_{2}}{r}\right)^{s}+sb_{2}\left(1-\frac{b_{2}}{r}\right)^{s-1}}{r^{4}\left(1-\frac{b_{2}}{r}\right)^{2s}}\right]r^{2}\left(1-\frac{b_{2}}{r}\right)^{s}}$ (57)
+$-\left[\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{p}+\frac{\beta Q^{2}}{r^{2}\left(1-\frac{b_{2}}{r}\right)^{s}}\right]\times\left[2r\left(1-\frac{b_{2}}{r}\right)^{s}+sb_{2}\left(1-\frac{b_{2}}{r}\right)^{s-1}\right].$
+This equation is highly nonlinear and cannot be solved analytically in closed form for arbitrary parameter values. Therefore, the photon sphere radius can only be expressed implicitly.
+However, once specific values of $\alpha,\beta,Q,M$ are chosen, the implicit equation can be solved numerically to yield $r_{\mathrm{ph}}$ with high precision using standard computational tools.
+This approach allows for direct investigation of the dependence of the photon sphere radius on the physical and coupling parameters of the EMS black hole.
+Appendix B Derivations for $\hat{\alpha}$
+We follow the Asahi Ishihara’s method 666 to calculate the deflection angle. At first, we have
+$dt^{2}=\gamma_{ij}dx^{i}dx^{j},$ (58)
+where $i$ and $j$ denote 1,2 and 3 , and $\gamma_{ij}$ is often called the optical metric. Then we can have the Gaussian Curvature
+$K=\frac{R_{r\phi r\phi}}{g},\quad g=\det(\gamma_{ij})=\frac{C(r)}{f(r)^{3}},$ (59)
+and the Gaussian Curvature then can be written as
+$K=-\frac{1}{2}\sqrt{\frac{f^{3}(r)}{C(r)}}\frac{\partial}{\partial r}\left(\sqrt{\frac{f^{3}(r)}{C(r)}}\frac{\partial}{\partial r}\left(\frac{C(r)}{f(r)}\right)\right),$ (60)
+thus (in the weak limit $M\ll r$)
+$K\approx$ $\frac{1}{r^{6}\left(r^{2}-Mrq^{2}\alpha^{2}\right)^{2}}\left(M(3M-2r)r^{6}+Mr^{2}rq^{2}\left(r^{4}+3M^{3}rq^{2}-2M^{2}r\left(r+rq^{2}\right)\right)\alpha^{2}\right$. (61)
+$+Q^{2}(2M-r)r\left(3r^{4}-4Mr^{2}rq^{2}\alpha^{2}+3M^{2}rq^{4}\alpha^{4}\right)\left(-1+\alpha^{2}-\beta\right)$
+$\left.+Q^{4}\left(2r^{4}-3Mr^{2}rq^{2}\alpha^{2}+2M^{2}rq^{4}\alpha^{4}\right)\left(1-\alpha^{2}+\beta\right)^{2}\right).$
+Next, we will use these equations to find the weak deflection angle using the Gauss-Bonnet theorem (GBT), originally stated as 667 ; 668
+$\iint_{D}KdS+\sum_{a=1}^{N}\int_{\partial D_{a}}\kappa_{\mathrm{g}}d\ell+\sum_{a=1}^{N}\theta_{a}=2\pi\chi(D).$ (62)
+[…]
+It was shown by 666 that in a SSS spacetime admitting asymptotic flatness, the deflection angle can be written as
+$\hat{\alpha}=-\iint_{D}K\,dS.$ (63)
+Taking the half-plane $\phi\in[0,\pi]$ with area element
+$dS\approx r+M\left(3-\frac{a^{2}q^{2}}{2}\right)+\frac{1}{r}\left[-\frac{3}{2}Q^{2}\left(1-a^{2}+\beta\right)+M^{2}\left(\frac{15}{2}-\frac{3a^{2}q^{2}}{2}-\frac{a^{4}q^{4}}{8}\right)\right]dr\,d\phi,$ (64)
+integrate over $r\in[r_{0},\infty]$.
+The final result is
+$\hat{\alpha}\approx\frac{4M}{b}-\frac{3\pi Q^{2}\left(1-\alpha^{2}+\beta\right)}{4b^{2}}.$ (65)